### v1

# Vector-Quantized Vision Foundation Models for Object-Centric Learning

Abstract
Decomposing visual scenes into objects, as humans do, facilitates modeling object relations and dynamics. Object-Centric Learning (OCL) achieves this by aggregating image or video feature maps into object-level feature vectors, known as slots. OCL’s self-supervision via reconstructing the input from slots struggles with complex textures, thus many methods employ Vision Foundation Models (VFMs) to extract feature maps with better objectness. However, using VFMs merely as feature extractors does not fully unlock their potential. We propose Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO), where VFM features are extracted to facilitate object-level information aggregation and further quantized to strengthen supervision in reconstruction. Our VVO unifies OCL representatives into a concise architecture. Experiments demonstrate that VVO not only outperforms mainstream methods on object discovery tasks but also benefits downstream tasks like visual prediction and reasoning. The source code is available in the supplement.
Object-Centric Learning, Vision Foundation Models, Vector Quantization, Object Discovery, Visual Prediction, Visual Reasoning
Aalto University, Espoo, Finland
{rongzhen.zhao, vivienne.wang, juho.kannala, joni.pajarinen}@aalto.fi
DINO2 ViT (s/14) ResNet18 (OCL fine-tune) distribution shift
intra-obj dist inter-obj dist intra-obj dist inter-obj dist (in PCA space)
COCO 0.6810 0.8772 0.8028 0.7984 0.7358
Figure 1: We utilize two observations: (i) Compared to non-VFMs, VFMs extract features with better object separability, i.e., smaller intra-object distances and larger inter-object distances $\Rightarrow$ We facilitate OCL aggregation via VFM features; (ii) VFM and non-VFM features have a distribution gap, i.e., far-flung centroids $\Rightarrow$ We strengthen OCL supervision by sharing the aforementioned VFM features and quantizing them as the reconstruction target.
1 Introduction
Objects can form highly diverse visual scenes through arrangements and combinations. But mainstream methods based on feature patches or a single feature vector disregard such compositionality. Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri & Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects. Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a).
Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020). Firstly, the encoder transforms input image or video frame pixels into a dense feature map. Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots. Lastly, the decoder reconstructs the input in some form from these slots, to provide the self-supervised training signal.
Due to its self-supervision nature, OCL heavily relies on pixel textures to separate different objects. The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects with complex textures. Some methods (Kipf et al., 2022; Elsayed et al., 2022) reconstruct the optical flow or depth map of the input to mitigate textural noises, albeit at the cost of extra expensive annotations. Some others (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots. Recent advances (Seitzer et al., 2023; Wu et al., 2023b) leverage Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract feature maps with better object separability, significantly improving OCL performance.
However, these methods merely treat VFMs as stronger feature extractors for OCL encoding, overlooking their potential to provide better reconstruction targets for OCL decoding. VFMs have yet to be fully utilized in OCL.
We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, VQ-VFM-OCL or VVO. As shown in Figure 1, we fully utilize VFMs in OCL in two ways. We not only use VFMs to extract feature maps with better objectness, to facilitate object information aggregation. We also quantize VFM feature maps with codebook codes, to strengthen supervision in reconstruction. Such shared quantization suppresses redundant feature information and produces consistent reconstruction targets across samples, facilitating aggregation via those initial slots, which are also shared across samples.
Our contributions are: (i) A unified architecture of encoder-aggregator/quantizer-decoder, which supports various OCL modules; (ii) Three key designs of direct VFM feature extraction, shared VFM feature quantization to fully utilize VFMs in OCL, and a VQ variant designed for OCL; (iii) Significant improvements in mainstream OCL methods, along with benefits to downstream task performance.
2 Related Work
SSseg vs OCL. Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024) and OCL can discovery objects, i.e., segment object, in images or videos without supervision. But SSseg only outputs object segmentation masks, while OCL represents objects as slots with segmentation as pyproduct. Besides, these slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks, like prediction, reasoning, planning and decision-making. A direct comparison between OCL and SSseg methods is beyond the scope of this work.
OCL encoding. The stronger encoder that extracts feature maps of better objectness contributes more to OCL performance. Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder. Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them in OCL. State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly. Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also recognized VFMs, they remain unexploited in OCL setting. Our VVO supports various VFMs for OCL encoding.
OCL aggregation. SlotAttention (Locatello et al., 2020) is the footstone for mainstream OCL methods. Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are all designed without changing the external interface. But considering their performance boosts, where some are very limited, we only integrate BO-QSA by default.
OCL decoding. With SlotAttention as the aggregator, the decoder and its reconstruction target affect OCL performance the most, as it is the source of supervision. Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (Zadaianchuk et al., 2024), decodes each slot’s spatial broadcast (Watters et al., 2019) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction. Transformer-based decoding, used in SLATE, STEVE and SPOT (Kakogeorgiou et al., 2024), reconstructs VAE representation of the input auto-regressively with slots as the condition. Diffusion-based decoding, used in LSD (Jiang et al., 2023) and SlotDiffusion, drives slots to recover noise added to the input’s VAE representation. Our VVO supports all these types of OCL decoding.
VAE for OCL. Variational Autoencoders (VAEs), like dVAE (Im Im et al., 2017) in SLATE and VQ-VAE (Van Den Oord et al., 2017) in SlotDiffusion, are employed to produce reconstruction targets for OCL training. Since these VAEs are designed for image generation, some methods adapt them for OCL. Inspired by channel or weight grouping, GDR (Zhao et al., 2024a) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better. MSF (Zhao et al., 2024b) exploits the multi-scale idea from detection and segmentation in OCL with novel VAE designs. Based on recent advancement RVQ (Yang et al., 2023) and SimVQ (Zhu et al., 2024), we design our own VQ variant for OCL.
Figure 2: VVO is a unified architecture that fully utilizes VFMs in OCL. It not only extracts VFM features with better objectness to facilitate object information aggregation; but further quantizes those VFM features as reconstruction targets to strengthen OCL training supervision. With typical SlotAttention or it variants as the aggregator and Vector-Quantization as the quantizer, VVO supports different VFMs as the encoder, and supports mainstream mixture, auto-regression and diffusion models as the decoder.
3 Proposed Method
We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, or VQ-VFM-OCL (VVO), unifying mainstream OCL with significant performance boosts. Using VFMs for better performance is not new, but VVO is the first to fully utilize VFMs in the OCL setting.
3.1 Unify OCL
As shown in Figure 2, our method adopts an architectural design of encoder-aggregator/quantizer-decoder.
Firstly, OCL encoder $\bm{\phi}_{\mathrm{e}}$ transforms an input image or video frame $\bm{X}\in\mathbb{R}^{h_{0}\times w_{0}\times c_{0}}$, which is an observation of some visual scene, into a dense feature map $\bm{Z}\in\mathbb{R}^{h\times w\times c}$, for the following query-based aggregation:
$\bm{\phi}_{\mathrm{e}}:\bm{X}\rightarrow\bm{Z}$ (1)
where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained frozen VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024). As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here. We will detail the reasons in Section 3.2.
Secondly, given queries $\bm{S}_{0}\in\mathbb{R}^{n\times c}$, OCL aggregator $\bm{\phi}_{\mathrm{a}}$ transforms $\bm{Z}$ into multiple feature vectors or slots $\bm{S}\in\mathbb{R}^{n\times c}$ and byproduct segmentation masks $\bm{M}\in\mathbb{R}^{h\times w}$, corresponding to objects or the background in the scene:
$\bm{\phi}_{\mathrm{a}}:(\bm{S}_{0},\bm{Z})\rightarrow(\bm{S},\bm{M})$ (2)
where $\bm{\phi}_{\mathrm{a}}$ can be parameterized as widely adopted SlotAttention (Locatello et al., 2020) and its variants, which is some cross attention with $\bm{S}_{0}$ as queries and $\bm{Z}$ as keys and values. $\bm{M}$ is the binarized attention map thus intuitively reflects how well objects are represented by slots.
For video OCL, there is a recurrent module turning current $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step. Such module can be Transformer encoder blocks (Vaswani et al., 2017).
Meanwhile, OCL quantizer $\bm{\phi}_{\mathrm{q}}$ transforms $\bm{X}$ into the reconstruction target $\bm{Q}\in\mathbb{R}^{h\times w\times c}$ for the following decoding:
$\bm{\phi}_{\mathrm{q}}:\bm{X}\rightarrow\bm{Q}$ (3)
where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) module (Im Im et al., 2017; Van Den Oord et al., 2017). But we meticulously design our own VQ variant, as detailed in Section 3.2. $\bm{\phi}_{\mathrm{q}}$ is pretrained in VAE framework and is frozen afterwards, where the encoder is shared from the frozen OCL encoder $\bm{\phi}_{\mathrm{e}}$, and the decoder is a typical VAE decoder. We will explain why not use a separate typical VAE encoder in Section 3.2.
Thirdly, OCL decoder $\bm{\phi}_{\mathrm{d}}$ transforms $\bm{S}$ into reconstruction $\bm{Q}^{\prime}\in\mathbb{R}^{h\times w\times c}$ with destructed $\bm{Q}$ as the condition:
$\bm{\phi}_{\mathrm{d}}:(\bm{S},\bm{Q})\rightarrow\bm{Q}^{\prime}$ (4)
where $\bm{\phi}_{\mathrm{d}}$ can be parameterized as a CNN or MLP if using mixture decoding for OCL (Kipf et al., 2022; Seitzer et al., 2023), where each of $\bm{S}$ is spatially broadcast (Watters et al., 2019) into the shape of $\bm{Q}$ and then decoded into a component and finally mixed together with predicted weights; or (ii) a Transformer decoder if using auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where causally masked $\bm{Q}$ is the query and $\bm{S}$ is the key and value; or (iii) a conditional Diffusion model if using diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ with noise is the input and $\bm{S}$ is the condition.
Reconstructing the destructed $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible. Thus, a good reconstruction target, like in (Zhao et al., 2024a, b), is very important. We will explain this in Section 3.2.
Lastly, the supervision signal for OCL training comes from minimizing the reconstruction loss between $\bm{Q}^{\prime}$ and $\bm{Q}$:
$\mathrm{min}_{(\bm{\phi}_{\mathrm{a}},\bm{\phi}_{\mathrm{d}})}f_{\mathrm{recon}}(\bm{Q}^{\prime},\bm{Q})$ (5)
where $f_{\mathrm{recon}}(\cdot,\cdot)$ can be (i) Mean Squared Error (MSE) loss for mixture and diffusion OCL decoding, or (ii) Cross Entropy (CE) loss for auto-regressive decoding.
3.2 Fully Utilize VFMs in OCL
We fully utilize VFMs in the OCL setting by two key designs: (i) direct VFM feature extraction to facilitate object-level information aggregation, and (ii) shared VFM feature quantization as reconstruction targets to strengthen supervision in OCL training.
Figure 3: Better objectness helps OCL aggregation. Green and orange areas stand for objects $\bm{o}_{1}$ and $\bm{o}_{2}$, where $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary super-pixels and $\bm{s}_{*}$ is $\bm{o}_{2}$’s centroid. But the actual query $\bm{s}\sim\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2})$. In VFM super-pixel space, the distance $d_{12}$ between $\bm{v}_{1}$ and $\bm{v}_{2}$ is larger, i.e., better objectness, thus $\bm{s}$ has higher probability $p(d_{s1}>d_{s2})$ to represent $\bm{o}_{2}$ correctly, compared with that in non-VFM super-pixel space.
Direct VFM Feature Extraction for Better Aggregation
We directly extract the feature map $\bm{Z}$ from the input $\bm{X}$ using VFMs as $\bm{\phi}_{\mathrm{e}}$, where DINO2 ViT and SAM2 encoder are chosen and experimented thoroughly. No extra position encoding is needed here like in SA (Locatello et al., 2020) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$. Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly. Thus, We expect the subsequent $\bm{\phi}_{\mathrm{a}}$ to aggregate dense $\bm{Z}$ into sparse slots $\bm{S}$ under queries $\bm{S}_{0}$ better.
Direct use of VFMs for feature extraction contributes to better aggregation. As shown in Figure 1, VFM representations show better objectness or object separability than that of non-VFMs, under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.clust
er.KMeans.html
. OCL aggregation based on SlotAttention and its variants is essentially a clustering whose initial centroids are trainable (Jia et al., 2023). This has already been empirically exploited in previous methods like DINOSAUR (Seitzer et al., 2023) and SlotDiffusion (Wu et al., 2023b), but without any explanation. We provide math analysis below.
Math Analysis: VFM Representation Eases Aggregation
As shown in Figure 3, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).
Denote the ideal query of $\bm{o}_{2}$ as $\bm{s}_{*}$, which is the clustering center or centroid of $\bm{o}_{2}$ and is closer to all super-pixels in $\bm{o}_{2}$ than in $\bm{o}_{1}$:
$d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$ (6)
where $d(\cdot,\cdot)$ is a distance metric, e.g., minus inner product; $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary points in $\bm{o}_{1}$ and $\bm{o}_{2}$, respectively.
But the actual query $\bm{s}$ follows $\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2}\bm{I})$. Substituting $\bm{s}$ for $\bm{s}_{*}$, the probability of correct aggregation is:
[…]
where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation plane between $\bm{o}_{1}$ and $\bm{o}_{2}$. The closer the boundary is to $\bm{s}_{*}$ the smaller $p_{2}$ is.
According to Figure 1, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space. This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space. Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space. $\square$
Figure 4: Qualitative object discovery performance comparison.
Shared VFM Feature Quantization for Better Supervision
Given the direct VFM feature map $\bm{Z}$, we transform it with a naive CNN to eliminate the positional information for better quantization. We quantize the adjusted $\bm{Z}$ using our VQ variant as $\bm{\phi}_{\mathrm{q}}$ to obtain the reconstruction target $\bm{Q}$.
Our VQ’s codebook follows SimVQ (Zhu et al., 2024). We predefine $m=4096$ template features, i.e., a codebook $\bm{T}_{0}\in\mathbb{R}^{m\times c_{0}}$, which are randomly initialized and remain frozen. In vector quantization, we project $\bm{T}$ with a pre-trainable linear layer and match it with the adjusted $\bm{Z}$:
$\bm{T}:=\bm{W}\cdot\mathrm{sg}(\bm{T}_{0})$ (8)
$\bm{D}=||\bm{Z}-\bm{T}||_{2}^{2}$ (9)
where $\mathrm{sg}(\cdot)$ is stop-gradient; $\bm{T}\in\mathbb{R}^{m\times c}$ is the codebook to be used for quantizing $\bm{Z}$; $\bm{D}\in\mathbb{R}^{h\times w\times m}$ is the matching distance between every super-pixel and every code.
We convert distances to probabilities and select the most matched codes to form the quantization $\bm{Q}$:
$\bm{P}=\mathrm{softmax}_{c}(-\bm{D})$ (10)
$\bm{I}=\mathrm{argmax}_{m}(\bm{P})$ (11)
$\bm{Q}=\mathrm{index}_{m}(\bm{T},\bm{I})$ (12)
where $\mathrm{softmax}(\cdot)$ is calculated along the channel dimension; $\bm{P}$ is the match probabilities; $\bm{I}\in\mathbb{R}^{h\times w}$ is the matched code indexes; $\mathrm{argmax}(\cdot)$ is calculated along the channel dimension; $\mathrm{index}(\cdot,\cdot)$ is operated along code number dimension. The typical STE (Bengio et al., 2013) on $\bm{Q}$, needed in pre-training, can be skipped during OCL training.
For its pre-training, we introduce some tricks. We add noise to $\bm{D}$ pre-Equation 10 to encourage code utilization:
$\bm{D}:=\frac{\bm{D}+\bm{G}}{\tau}$ (13)
where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is the Gumbel noise and $\tau$ is the temperature. We devise annealing residual connection (Zhao et al., 2024a, b) post-Equation 12 to stabilize pre-training:
$\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$ (14)
where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing. Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$’s elements to be normalized:
$l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$ (15)
where $\lambda$ is empirically set to 0.1; $\mathbb{E}$ and $\mathbb{V}$ are calculated along height, width and channel dimensions.
With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency and suppresses information redundancy or noise. This helps the aggregation with queries $\bm{S}_{0}$, which are also shared across samples. After such tokenization, our reconstruction target can be either regressed using mixture and diffusion decoding or classified using auto-regressive decoding. In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) use separate VAE and OCL encoders, which as shown in Figure 1 causes a distribution gap between $\bm{Q}$ and $\bm{Z}$. Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen supervision.
Math Analysis: Shared VFM Representation Quantization Reduces Supervision Noise
We reconstruct $\bm{Q}^{\prime}$ to approximate the target $\bm{Q}$ ultimately from $\bm{Z}$ via $\bm{\phi}_{\mathrm{a}}\circ\bm{\phi}_{\mathrm{d}}$, denoted as $f$ for simplicity. Under MSE loss, the gradient with respect to $\bm{Z}$ is:
$\frac{\partial\mathrm{MSE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=2(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))\frac{\partial\bm{Q}^{\prime}}{\partial\bm{Z}}$ (16)
We obtain $\bm{Q}$ by quantizing $\bm{Z}$, i.e., $\mathbb{E}[\bm{Z}]=\bm{Q}$, implying that any deviation of $\bm{Q}^{\prime}=f(\bm{Z})$ from $\bm{Q}$ is due to $f$ and $\bm{Z}$. Assuming $f$ preserves $\bm{Z}$’s statistical properties, we have:
$\mathbb{E}[\bm{Q}^{\prime}]=\mathbb{E}[f(\bm{Z})]\approx\bm{Q}$ (17)
Thus the residual error $\bm{Q}^{\prime}-\mathrm{sg}({\bm{Q}})$ in Equation 16 is statistically unbiased and small on average:
$\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q})]\approx 0$ (18)
But if instead of sharing $\bm{\phi}_{\mathrm{e}}$, we use an extra VAE encoder plus $\bm{\phi}_{\mathrm{a}}$ to obtain the target, denoted as $\bm{Q}_{2}$, then $\bm{Q}_{2}\neq\bm{Q}$. Substitute $\bm{Q}_{2}$ into Equation 16 and the residual error $\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})$ would be systematically biased:
$\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})]\neq 0$ (19)
which increases noise in optimization.
Using CE loss is slightly different but it does not alter our judgment above. $\square$
resolution= ClevrTex #slot=11 COCO (panoptic) #slot=7 VOC #slot=6 MOVi-D #slot=21, conditional
256$\times$256 ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO
SLATE-DINO 15.8
$\pm$0.3
85.8
$\pm$0.3
26.2
$\pm$0.1
40.0
$\pm$0.2
34.4
$\pm$0.3
35.4
$\pm$0.5
7.1
$\pm$0.1
26.8
$\pm$0.7
16.4
$\pm$0.6
22.2
$\pm$0.9
13.0
$\pm$0.4
34.7
$\pm$0.9
VQDINO
Tfd
21.5 $\pm$0.2 86.1 $\pm$0.1 29.0 $\pm$0.1 43.3 $\pm$0.1 38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4 19.0 $\pm$0.9 24.1 $\pm$0.7 14.3 $\pm$0.3 37.2 $\pm$0.8
STEVE-DINO 35.0
$\pm$0.9
64.1
$\pm$2.9
21.2
$\pm$0.7
23.2
$\pm$1.2
VQDINO
Tfdt
41.4 $\pm$0.8 70.5 $\pm$3.3 23.1 $\pm$0.6 25.9 $\pm$1.4
DINOSAUR-DINO 18.5
$\pm$0.2
87.8
$\pm$0.4
26.7
$\pm$0.3
39.9
$\pm$0.5
35.3
$\pm$0.1
36.4
$\pm$0.2
7.1
$\pm$0.0
27.3
$\pm$0.2
18.8
$\pm$0.4
31.2
$\pm$0.1
14.4
$\pm$0.2
37.9
$\pm$0.4
VQDINO
MLP
51.2 $\pm$0.4 88.3 $\pm$0.3 29.8 $\pm$0.2 44.6 $\pm$0.4 35.4 $\pm$0.6 36.7 $\pm$0.5 7.6 $\pm$0.1 27.7 $\pm$0.1 20.6 $\pm$0.1 32.9 $\pm$0.1 15.6 $\pm$0.0 39.1 $\pm$0.3
SlotDiffusion-DINO 60.1
$\pm$9.4
81.6
$\pm$2.7
31.4
$\pm$0.9
45.3
$\pm$0.9
34.9
$\pm$0.4
36.0
$\pm$0.4
7.0
$\pm$0.2
26.7
$\pm$0.6
15.0
$\pm$0.3
19.7
$\pm$0.4
12.3
$\pm$0.2
33.1
$\pm$0.3
VQDINO
Dfz
37.9 $\pm$2.1 81.7 $\pm$2.6 29.3 $\pm$1.1 47.8 $\pm$1.4 37.2 $\pm$0.7 38.3 $\pm$0.9 7.3 $\pm$0.1 27.8 $\pm$0.9 17.8 $\pm$0.1 24.1 $\pm$0.3 14.1 $\pm$0.1 36.1 $\pm$0.3
Table 1: Object discovery performance with DINO2 ViT (s/14) for OCL encoding. VVO is instantiated as VQDINO; Tfd, Tfdt, MLP and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
resolution= ClevrTex #slot=11 COCO (panoptic) #slot=7 VOC #slot=6 MOVi-D #slot=21, conditional
256$\times$256 ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO
SLATE-SAM 12.4
$\pm$1.1
80.3
$\pm$0.6
28.4
$\pm$0.3
44.0
$\pm$0.8
34.7
$\pm$0.6
35.8
$\pm$1.4
6.4
$\pm$0.3
25.9
$\pm$0.4
12.7
$\pm$0.4
19.5
$\pm$0.9
11.7
$\pm$0.2
31.6
$\pm$1.1
VQSAM
Tfd
14.2 $\pm$1.6 81.1 $\pm$0.2 28.5 $\pm$0.4 44.8 $\pm$0.6 36.5 $\pm$0.8 37.9 $\pm$0.8 7.4 $\pm$0.1 27.9 $\pm$0.5 17.3 $\pm$0.6 19.6 $\pm$0.7 13.1 $\pm$0.2 34.5 $\pm$0.8
STEVE-SAM 28.0
$\pm$0.5
59.3
$\pm$1.2
18.0
$\pm$0.3
20.3
$\pm$0.9
VQSAM
Ttdt
39.7 $\pm$0.2 68.5 $\pm$0.8 24.2 $\pm$0.2 26.3 $\pm$0.4
DINOSAUR-SAM 14.4
$\pm$0.3
87.6
$\pm$0.9
25.8
$\pm$0.7
39.3
$\pm$0.7
27.3
$\pm$0.4
27.7
$\pm$0.5
5.9
$\pm$0.1
22.6
$\pm$0.5
11.8
$\pm$0.7
16.7
$\pm$1.1
11.2
$\pm$0.3
29.6
$\pm$1.2
VQSAM
MLP
25.3 $\pm$2.6 84.1 $\pm$0.2 29.4 $\pm$1.1 44.1 $\pm$0.9 36.5 $\pm$0.8 37.6 $\pm$0.8 7.2 $\pm$0.1 27.6 $\pm$0.3 17.3 $\pm$0.6 19.2 $\pm$0.7 12.8 $\pm$0.4 32.2 $\pm$0.8
SlotDiffusion-SAM 19.0
$\pm$1.0
76.7
$\pm$2.3
31.2
$\pm$0.5
47.3
$\pm$1.2
34.5
$\pm$0.3
35.6
$\pm$0.3
7.0
$\pm$0.1
26.7
$\pm$0.3
15.2
$\pm$0.6
17.2
$\pm$1.3
12.6
$\pm$0.7
33.9
$\pm$1.6
VQSAM
Dfz
16.7 $\pm$0.3 76.9 $\pm$0.3 32.5 $\pm$0.7 49.2 $\pm$0.9 35.7 $\pm$0.6 36.8 $\pm$0.6 7.2 $\pm$0.0 27.1 $\pm$0.4 16.9 $\pm$1.3 19.6 $\pm$1.4 13.6 $\pm$0.2 36.0 $\pm$0.9
Table 2: Object discovery performance with SAM2 Hiera+FPN (t/16) for OCL decoding. VVO is instantiated as VQSAM; Tfd, Tfdt, MLP and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
4 Experiment
We conduct all experiments using three random seeds.
4.1 Set up the Benchmark
Datasets. We include both synthetic and real-world datasets. ClevrTex [^2] https://www.robots.ox.ac.uk/~vgg/data/clevrtex comprises synthetic images, each with about 10 geometric objects scattered in complex background. MOVi-D [^3] https://github.com/google-research/kubric/blob/main/challeng
es/movi/README.md#movi-d
contains synthetic videos, each with up to 20 daily objects dropping and bumping. COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its challenging panoptic segmentation. VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation. We also report results on real-world video dataset YTVIS [^6] https://youtube-vos.org/dataset/vis version HQ [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube. We choose Physion [^8] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
Models. We compare VVO with both OCL classics and state-of-the-arts. SLATE (Singh et al., 2022a) uses a Transformer decoder for auto-regressive decoding, and it differs from VVO in a separate VAE encoder and naive quantizer; STEVE (Singh et al., 2022c) is SLATE’s video version. DINOSAUR (Seitzer et al., 2023) uses an MLP for mixture decoding, and it differs from VVO in no quantization in its reconstruction target. SlotDiffusion (Wu et al., 2023b) uses a conditional Diffusion model for diffusion decoding, and it differs from VVO in a separate VAE encoder and naive quantizer. General improvers GDR (Zhao et al., 2024a) and MSF (Zhao et al., 2024b) only support auto-regression and diffusion decoding. We skip outdated methods like IODINE (Greff et al., 2019), SA (Locatello et al., 2020) and ISA (Biza et al., 2023) due to their low accuracy. We also skip SAVi (Kipf et al., 2022) and SAVi++ (Elsayed et al., 2022) as their extra modalities are unfair to others.
Comparison. Instead of copying existing results, we reproduce all baselines to realize fair comparison. As detailed in Appendices A and B, we use identical data augmentation, VFMs in OCL encoding and training recipes for all experiment items unless not applied. We instantiate all baselines’ VAE part as TAESD [^9] https://huggingface.co/docs/diffusers/en/api/models/autoenco
der_tiny
, which is a large-scale pretrained StableDiffusion [^10] https://huggingface.co/spaces/stabilityai/stable-diffusion module, to build all strong baselines.
ARI ARIfg mIoU mBO
Using higher resolution: 384$\times$384
resolution=384$\times$384 COCO (panoptic) #slot=7
SLATE-DINO 50.9
$\pm$0.3
51.0
$\pm$0.9
9.0
$\pm$0.1
27.9
$\pm$0.7
VQDINO
Tfd
55.6 $\pm$0.4 55.9 $\pm$0.3 9.8 $\pm$0.0 30.2 $\pm$0.2
DINOSAUR-DINO 53.8
$\pm$0.2
54.1
$\pm$0.4
9.2
$\pm$0.0
28.9
$\pm$0.3
VQDINO
MLP
54.7 $\pm$0.3 55.2 $\pm$0.6 9.4 $\pm$0.1 29.8 $\pm$0.3
SlotDiffusion-DINO 53.1
$\pm$0.6
53.3
$\pm$0.8
9.1
$\pm$0.2
28.4
$\pm$0.7
VQDINO
Dfz
55.3 $\pm$0.8 55.7 $\pm$0.5 9.8 $\pm$0.1 30.2 $\pm$0.4
Using different aggregators: SlotAttention, BO-QSA
resolution=256$\times$256 COCO (panoptic) #slot=7
SLATE-DINO-SlotAttention 34.1
$\pm$0.7
34.9
$\pm$0.6
7.0
$\pm$0.1
26.2
$\pm$0.9
VQDINO Tfd -SlotAttention 37.7 $\pm$0.5 39.3 $\pm$0.4 7.7 $\pm$0.1 28.0 $\pm$0.6
SLATE-DINO-BO-QSA 34.4
$\pm$0.3
35.4
$\pm$0.5
7.1
$\pm$0.1
26.8
$\pm$0.7
VQDINO Tfd -BO-QSA 38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4
Table 3: VVO using higher resolution (upper) and different aggregators (lower) on object discovery. By default, we use BO-QSA for all our experiment items, including the baselines.
ARI ARIfg mIoU mBO
Compared with general improvers: GDR and MSF
resolution=256$\times$256 COCO (panoptic) #slot=7
GDR
Tfd
-DINO
34.8
$\pm$0.4
36.0
$\pm$0.6
7.1
$\pm$0.1
27.1
$\pm$0.7
MSF
Tfd
-DINO
35.9
$\pm$0.3
37.1
$\pm$0.4
7.2
$\pm$0.0
27.4
$\pm$0.5
VQDINO
Tfd
38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4
GDR
Dfz
-DINO
36.2
$\pm$0.5
37.1
$\pm$0.4
7.1
$\pm$0.1
27.2
$\pm$0.6
MSF
Dfz
-DINO
35.6
$\pm$0.2
36.6
$\pm$0.4
7.1
$\pm$0.0
27.2
$\pm$0.4
VQDINO
Dfz
37.2 $\pm$0.7 38.3 $\pm$0.9 7.3 $\pm$0.1 27.8 $\pm$0.9
Compared with SotA methods: SPOT, VideoSAUR
resolution=256$\times$256 COCO (panoptic) #slot=7
SPOT-DINO 37.5
$\pm$0.4
38.7
$\pm$0.3
7.8
$\pm$0.1
28.1
$\pm$0.9
VQDINO
Tfd9
+self-training
39.3 $\pm$0.3 40.4 $\pm$0.4 8.0 $\pm$0.0 29.6 $\pm$0.5
resolution=256$\times$256 YTVIS (HQ) #slot=7, unconditional
VideoSAUR-DINO 34.0
$\pm$0.6
39.1
$\pm$0.9
27.2
$\pm$0.4
32.2
$\pm$0.6
VQDINO
Smd
40.7 $\pm$0.5 43.5 $\pm$0.6 29.2 $\pm$0.2 35.1 $\pm$0.5
Table 4: VVO compared with general improvers (upper) and SotA methods (lower) on object discovery. SPOT uses both Transformer decoder with 9 permutations Tfd9 and “self-training”; VideoSAUR uses SlotMixer Smd as its decoder.
4.2 Evaluate on Object Discovery
Object discovery task intuitively shows how well those slots separate different objects. We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.adjusted_rand_score.html
(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.jaccard_score.html
(mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO. As shown in Table 1, our method consistently improves object discovery performance across all types of OCL decoding. With a Transformer decoder Tfd for auto-regressive decoding, VVO significantly outperforms SLATE and STEVE across all datasets. With MLP Mlp for mixture decoding, VVO shows a smaller advantage over DINOSAUR but is still effective. With a conditional Diffusion model Dfz for diffusion decoding, VVO surpasses SlotDiffusion on most datasets.
With supervised pretrained VFMs for OCL encoding, i.e., SAM2 Hiera+FPN (version t/16), our VVO is instantiated as VQSAM. As shown in Table 2, VVO consistently boosts those baseline methods’ object discovery performance across all decoding types on most datasets.
As shown in Table 3, whether using a higher input resolution or different aggregators, VVO maintains its superiority over baselines. As shown in Table 4, VVO outperforms recent general improvers for OCL training, i.e., GDR and MSF, which do not support mixture decoding like us though. VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
4.3 Evaluate on Set Prediction
class labels bounding boxes
top1$\uparrow$ R2$\uparrow$
DINOSAUR + MLP 0.61
$\pm$0.0
0.57
$\pm$0.1
VQDINO
MLP
+ MLP
0.64 $\pm$0.1 0.59 $\pm$0.1
Table 5: Set prediction performance on COCO (panoptic, #slot=7).
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023). We measure top1 accuracy of the classified class labels while we measure the R2 score [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.r2_score.html
of the regressed bounding box coordinates.
As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO-MLP, obtains better object classification and bounding box regression. Thus, our models’ object representations are of higher quality.
4.4 Deploy to the Downstream
Figure 5: Visual prediction (upper) and reasoning (lower) performance on Physion (#slot=8). VVO has smaller prediction error in all time steps, and higher reasoning accuracy in later time steps.
Better object representation benefits downstream tasks. We follow the convention to pretrain OCL models on Physion and represent this dataset as slots. Then the object-centric dynamics model SlotFormer (Wu et al., 2023a) is trained on those slots in an auto-regressive manner along the time dimension. We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO-MLP to extract slots.
On visual prediction, we evaluate the prediction errors per time step measured in normalized MSE between regressed and extracted slots. As shown in Figure 5 upper, VVO accumulates prediction errors slower than the baseline.
On visual reasoning, we evaluate the reasoning accuracies per time step between classified and ground-truth labels. As shown in Figure 5 lower, VVO accuracies are slightly lower at the beginning but much higher later than the baseline.
4.5 Ablate the Architecture
As shown in Table 6, VVO’s design of shared VAE and OCL encoder consistently outperforms separate VAE and OCL encoders, even when the latter employs another VFM for VAE encoding. Thus, VVO’s design of shared VFM representation quantization is crucial. However, the design of separate encoders remains prevalent among those baselines.
As shown in Table 7, VVO’s improved quantizer variant, based on tricks of Gumbel noises defined in Equation 13, annealing residual connection defined in Equation 14 and normalizing regularization defined in Equation 15, is superior to the naive VQ. The detailed effects of those tricks are shown in Figure 6. Adding Gumbel noises increases codebook utilization, contributing to more effective codes; Annealing residual connection improves VAE pretraining, contributing to smaller VAE reconstruction error.
VAE encoding OCL encoding VQDINO
Tfd
ARI ARIfg
shared DINO2 ViT DINO2 ViT 38.5
$\pm$0.4
39.6
$\pm$0.4
separate TAESD encoder ResNet18 28.3 $\pm$1.3 29.1 $\pm$1.5
TAESD encoder DINO2 ViT 34.6
$\pm$0.7
35.8
$\pm$0.5
SAM2 encoder DINO2 ViT 34.9
$\pm$0.3
35.8
$\pm$0.4
Table 6: VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders. Results are on COCO panoptic.
shared VAE / Gumbel annealing normaliz. VQDINO
Dfz
OCL encoder noise residual regulariz. ARI ARIfg
37.2
$\pm$0.7
38.3
$\pm$0.9
35.4
$\pm$2.5
36.6
$\pm$2.2
36.5
$\pm$1.6
37.6
$\pm$1.9
36.8
$\pm$1.3
37.8
$\pm$1.0
36.9
$\pm$0.7
37.9
$\pm$0.8
Table 7: VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts. In comparison to VVO’s key designs, these tricks are more like the cherry on top. Results are on COCO panoptic with settings consistent with the above.
Figure 6: Effects of tricks in our VQ variant: (upper) Gumbel noise improves codebook utilization, where “CV” means Coefficient of Variation, and curves are smoothed by Gaussian kernel of size 10; (lower) Annealing residual connection improves VAE pretraining, and the blue curve’s turning point at epoch 4 is where the residual connection anneals to zero. Results are from the VAE pretraining of VQDINO-Dfz on COCO panoptic.
Conclusion
We propose a unified architecture VVO for object-centric representation learning. Our VVO supports different well-recognized vision foundation models for OCL encoding and supports mainstream types of OCL decoding. It boosts the existing OCL performance in object discovery significantly, and benefits downstream tasks of visual prediction and reasoning. VVO has the potential to serve as a general testbed for research related to OCL in the future.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Dataset Processing
The datasets for image OCL are ClevrTex, COCO (panoptic) and VOC. The datasets for video OCL are MOVi-D and YTVIS. For different datasets, we have mostly the same processing but a bit differences.
A.1 Shared Preprocessing
To accelerate experiments, we adopt the dataset conversion scheme: Convert all datasets into LMDB database format, and load it into NVMe disk or RAM disk to minimize I/O overhead and maximize throughput.
Specifically, all images or video frames are resized by short side to 256, as well as their segmentation masks, and then are stored data type uint8 to save space. For images or videos, we use default interpolation mode; but for segmentation masks, we use NEAREST-EXACT interpolation mode.
During training, for both images and videos, we use spatial resolution of 256$\times$256 as inputs and outputs. We normalize the input images or videos by ImageNet-Mean and -Std, which are the common normalization scheme using by VFMs like DINO and SAM. For videos, we also using random strided temporal cropping of fixed window size 6 to accelerate training and improve generalization.
During testing, there is no difference to that of training for images. But for videos, we need to remove the random strided temporal cropping and use the full video of 24 time steps.
A.2 Dataset-Specific Processing
ClevrTex. As there are 10 objects at most in each image, we use number of slots queries of 10 + 1, where 1 represents for the background.
COCO & VOC. We use the panoptic segmentation annotations of COCO for evaluation, which is more challenging than instance segmentation. Besides, OCL’s queries-based workflow intuitively fits panoptic segmentation rather than instance segmentation or semantic segmentation. There can be lots of or only two objects in an image, but we just follow the popular hyperparameter selection, i.e., 7 slots for COCO and 6 slots for VOC, respectively.
MOVi-D. There are abundant annotations like depth, optical-flow, camera intrinsic/extrinsic parameters and so on. But for simple comparison, we only utilize the segmentation masks for evaluation and the bounding boxes as slots initialization. As there are 23 objects in these dataset, we use number of slots queries of 23+1.
YTVIS. We use the original YTVIS videos and HQ-YTVIS segmentation annotations. The latter provides high-quality masks. During VAE pretraining, we utilize all the video frames. But during OCL training, we only utilize the every-5-time-step annotations for calculating the performance.
Appendix B Training Recipe
Following the convention, there are two training stages. The pretraining stage is for training VAE modules on corresponding datasets, so as to learn good discrete intermediate representation. The OCL training stage, which utilize the pretraining VAE representation as guidance for object centric learning.
B.1 Pretrain VAE
For all datasets, we use total number of training iterations 30,000, and validation interval iterations 600, so that we have about 50 checkpoints for every OCL model on every dataset. But to save storage, we save only the latter half 25 checkpoints. This is unified across all datasets.
The batch size for image datasets is 64, while for video datasets it is 16. This holds for both training and validation. This setting is shared across all datasets.
For multi-process, we use the number of workers 4. This holds for both training and validation. This setting is shared across all datasets.
We use Adam optimizer with initial learning rate of 2e-3. And the learning rate is manipulated by cosine annealing scheduling, with a linear warmup during first 1/20 of total steps. This setting is shared across all datasets.
We use automatic mixed precision provided by PyTorch autocast API plus compile. Along with this, we use PyTorch builtin gradient scaler, so as to apply gradient clipping of maximum norm 1.0. This setting is shared across all datasets.
B.2 Train OCL Model
For all datasets, we use total number of training iterations 5,$0$, and validation interval iterations 1000, so that we have about 50 checkpoints for every OCL model on every dataset. But to save storage, we only save the latter half of the total checkpoints. This setting is shared across all datasets.
The batch size for image datasets is 32 for both training and validation, while for video datasets it is 8 for training and 4 for validation as there are more time steps in videos during validation. This setting is shared across all datasets.
For multi-process, we use the number of workers 4. This holds for both training and validation. This setting is shared across all datasets.
The objective is the CE and MSE loss following the basis methods original design.
Metrics used here are ARI, mIoU and mBO, calculating the panoptic segmentation accuracy of both objects and the background. This setting is shared across all datasets.
We use Adam optimizer with initial learning rate of 2e-4. And the learning rate is manipulated by cosine annealing scheduling, with a linear warmup during first 1/20 of total steps. This setting is shared across all datasets.
We use auto mixed precision provided by PyTorch autocast API and compile. Along with this, we use PyTorch builtin gradient scaler, so as to apply gradient clipping of maximum norm 1.0 for images and 0.02 for videos.

### v3

@@
 Vector-Quantized Vision Foundation Models for Object-Centric Learning
-Abstract
-Decomposing visual scenes into objects, as humans do, facilitates modeling object relations and dynamics.
-Object-Centric Learning (OCL) achieves this by aggregating image or video feature maps into object-level feature vectors, known as slots.
-OCL’s self-supervision via reconstructing the input from slots struggles with complex textures, thus many methods employ Vision Foundation Models (VFMs) to extract feature maps with better objectness.
-However, using VFMs merely as feature extractors does not fully unlock their potential.
-We propose Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO), where VFM features are extracted to facilitate object-level information aggregation and further quantized to strengthen supervision in reconstruction.
-Our VVO unifies OCL representatives into a concise architecture.
-Experiments demonstrate that VVO not only outperforms mainstream methods on object discovery tasks but also benefits downstream tasks like visual prediction and reasoning.
-The source code is available in the supplement.
-Object-Centric Learning, Vision Foundation Models, Vector Quantization, Object Discovery, Visual Prediction, Visual Reasoning
-Aalto University, Espoo, Finland
-{rongzhen.zhao, vivienne.wang, juho.kannala, joni.pajarinen}@aalto.fi
+Abstract.
+Perceiving visual scenes as objects and background—like humans do—Object-Centric Learning (OCL) aggregates image or video feature maps into object-level feature vectors, termed slots.
+OCL’s self-supervision of reconstructing the input from these aggregated slots struggles with complex object textures, thus Vision Foundation Model (VFM) representations are used as the aggregation input and reconstruction target.
+However, existing methods leverage VFM representations in diverse ways and often fail to fully exploit their potential.
+In response, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods.
+The key to our unification is simple yet effective, just shared quantizing the same VFM representation as the reconstruction target.
+Through mathematical modeling and statistical verification, we further analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision.
+Experiments show that across different VFMs, aggregators and decoders, our VVO consistently outperforms baselines in object discovery and recognition, as well as downstream visual prediction and reasoning.
+The implementation and model checkpoints are available on https://github.com/Genera1Z/VQ-VFM-OCL.
+isbn: 978-1-4503-XXXX-X/2018/06submissionid: 7307copyright: none
 DINO2 ViT (s/14) ResNet18 (OCL fine-tune) distribution shift
@@
 COCO 0.6810 0.8772 0.8028 0.7984 0.7358
-Figure 1: We utilize two observations: (i) Compared to non-VFMs, VFMs extract features with better object separability, i.e., smaller intra-object distances and larger inter-object distances $\Rightarrow$ We facilitate OCL aggregation via VFM features; (ii) VFM and non-VFM features have a distribution gap, i.e., far-flung centroids $\Rightarrow$ We strengthen OCL supervision by sharing the aforementioned VFM features and quantizing them as the reconstruction target.
-1 Introduction
+Figure 1.
+We utilize two observations: (i) Compared to non-VFMs, VFMs extract features with better object separability, i.e., smaller intra-object distances and larger inter-object distances $\Rightarrow$ We facilitate OCL aggregation via VFM features; (ii) VFM and non-VFM features have a distribution gap, i.e., separated centroids $\Rightarrow$ We strengthen OCL supervision by reconstructing the quantized features shared from the same VFM rather than from another encoder.
+1. Introduction
 Objects can form highly diverse visual scenes through arrangements and combinations.
 But mainstream methods based on feature patches or a single feature vector disregard such compositionality.
-Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri & Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects.
-Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a).
+Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects.
+Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024b).
 Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020).
 Firstly, the encoder transforms input image or video frame pixels into a dense feature map.
-Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots.
-Lastly, the decoder reconstructs the input in some form from these slots, to provide the self-supervised training signal.
-Due to its self-supervision nature, OCL heavily relies on pixel textures to separate different objects.
-The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects with complex textures.
-Some methods (Kipf et al., 2022; Elsayed et al., 2022) reconstruct the optical flow or depth map of the input to mitigate textural noises, albeit at the cost of extra expensive annotations.
-Some others (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots.
+Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots as the query.
+Lastly, the decoder reconstructs the input in some form from these aggregated slots, to provide the self-supervised training signal.
+OCL relies on pixel textures to discover objects.
+The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects.
+Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, but at the cost of expensive annotations.
+Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots.
 Recent advances (Seitzer et al., 2023; Wu et al., 2023b) leverage Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract feature maps with better object separability, significantly improving OCL performance.
-However, these methods merely treat VFMs as stronger feature extractors for OCL encoding, overlooking their potential to provide better reconstruction targets for OCL decoding.
-VFMs have yet to be fully utilized in OCL.
-We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, VQ-VFM-OCL or VVO.
-As shown in Figure 1, we fully utilize VFMs in OCL in two ways.
-We not only use VFMs to extract feature maps with better objectness, to facilitate object information aggregation.
-We also quantize VFM feature maps with codebook codes, to strengthen supervision in reconstruction.
-Such shared quantization suppresses redundant feature information and produces consistent reconstruction targets across samples, facilitating aggregation via those initial slots, which are also shared across samples.
-Our contributions are: (i) A unified architecture of encoder-aggregator/quantizer-decoder, which supports various OCL modules; (ii) Three key designs of direct VFM feature extraction, shared VFM feature quantization to fully utilize VFMs in OCL, and a VQ variant designed for OCL; (iii) Significant improvements in mainstream OCL methods, along with benefits to downstream task performance.
-2 Related Work
+However, existing OCL methods leverage VFM representations in quite different ways, as shown in Figure 3, and none of them fully utilize the power of VFM representations.
+To address these issues, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods.
+In this architecture, as shown in Figure 2, our VVO supports different VFMs for encoding, different OCL aggregators and different OCL decoders.
+The key to such unification is very simple—We quantize the representations from the same VFM, rather than from another encoder, as the reconstruction target.
+We also mathematically and statistically analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision.
+Our contributions are: (i) A clean architecture, which unifies mainstream OCL methods.
+(ii) Shared quantized VFM representations as reconstruction targets, which not only supports various OCL decoders but also boosts performance.
+(iii) Insights of why VFM features facilitate OCL aggregation and their shared quantization as reconstruction targets strengthens OCL supervision.
+2. Related Work
 SSseg vs OCL.
-Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024) and OCL can discovery objects, i.e., segment object, in images or videos without supervision.
-But SSseg only outputs object segmentation masks, while OCL represents objects as slots with segmentation as pyproduct.
-Besides, these slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks, like prediction, reasoning, planning and decision-making.
-A direct comparison between OCL and SSseg methods is beyond the scope of this work.
+Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024a) and OCL can discovery objects, i.e., segmenting objects and the background in images or video frames without supervision.
+SSseg only outputs object segmentation masks, while OCL further represents objects as slots.
+These slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks.
+Thus a direct comparison between them is beyond the scope of this work.
 OCL encoding.
@@
 Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder.
-Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them in OCL.
-State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly.
-Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also recognized VFMs, they remain unexploited in OCL setting.
+Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them on OCL datasets.
+State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly.
+Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs, they remain unexploited in OCL setting.
 Our VVO supports various VFMs for OCL encoding.
@@
 SlotAttention (Locatello et al., 2020) is the footstone for mainstream OCL methods.
-Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are all designed without changing the external interface.
-But considering their performance boosts, where some are very limited, we only integrate BO-QSA by default.
+Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are designed without changing the external interface.
+But considering their performance boosts, we only integrate BO-QSA by default.
 OCL decoding.
@@
 Transformer-based decoding, used in SLATE, STEVE and SPOT (Kakogeorgiou et al., 2024), reconstructs VAE representation of the input auto-regressively with slots as the condition.
-Diffusion-based decoding, used in LSD (Jiang et al., 2023) and SlotDiffusion, drives slots to recover noise added to the input’s VAE representation.
+Diffusion-based decoding in LSD (Jiang et al., 2023) and SlotDiffusion drives slots to recover noise added to the input’s VAE representation.
 Our VVO supports all these types of OCL decoding.
@@
 Inspired by channel or weight grouping, GDR (Zhao et al., 2024a) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better.
-MSF (Zhao et al., 2024b) exploits the multi-scale idea from detection and segmentation in OCL with novel VAE designs.
+MSF (Zhao et al., 2024b) firstly exploits the multi-scale idea in the OCL setting with VAE-specific designs.
 Based on recent advancement RVQ (Yang et al., 2023) and SimVQ (Zhu et al., 2024), we design our own VQ variant for OCL.
-Figure 2: VVO is a unified architecture that fully utilizes VFMs in OCL.
+Figure 2.
+VVO is a unified architecture that fully utilizes VFMs in OCL.
 It not only extracts VFM features with better objectness to facilitate object information aggregation; but further quantizes those VFM features as reconstruction targets to strengthen OCL training supervision.
 With typical SlotAttention or it variants as the aggregator and Vector-Quantization as the quantizer, VVO supports different VFMs as the encoder, and supports mainstream mixture, auto-regression and diffusion models as the decoder.
-3 Proposed Method
-We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, or VQ-VFM-OCL (VVO), unifying mainstream OCL with significant performance boosts.
-Using VFMs for better performance is not new, but VVO is the first to fully utilize VFMs in the OCL setting.
-3.1 Unify OCL
+3. Proposed Method
+We propose Vector-Quantized Vision Foundation Models for Object-Centric Learning, or VQ-VFM-OCL (VVO), elegantly unifying mainstream OCL and consistently boosting their performance.
+3.1. Unify OCL
 As shown in Figure 2, our method adopts an architectural design of encoder-aggregator/quantizer-decoder.
 Firstly, OCL encoder $\bm{\phi}_{\mathrm{e}}$ transforms an input image or video frame $\bm{X}\in\mathbb{R}^{h_{0}\times w_{0}\times c_{0}}$, which is an observation of some visual scene, into a dense feature map $\bm{Z}\in\mathbb{R}^{h\times w\times c}$, for the following query-based aggregation:
-$\bm{\phi}_{\mathrm{e}}:\bm{X}\rightarrow\bm{Z}$ (1)
+(1) $\bm{\phi}_{\mathrm{e}}:\bm{X}\rightarrow\bm{Z}$
 where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained frozen VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024).
@@
 Secondly, given queries $\bm{S}_{0}\in\mathbb{R}^{n\times c}$, OCL aggregator $\bm{\phi}_{\mathrm{a}}$ transforms $\bm{Z}$ into multiple feature vectors or slots $\bm{S}\in\mathbb{R}^{n\times c}$ and byproduct segmentation masks $\bm{M}\in\mathbb{R}^{h\times w}$, corresponding to objects or the background in the scene:
-$\bm{\phi}_{\mathrm{a}}:(\bm{S}_{0},\bm{Z})\rightarrow(\bm{S},\bm{M})$ (2)
+(2) $\bm{\phi}_{\mathrm{a}}:(\bm{S}_{0},\bm{Z})\rightarrow(\bm{S},\bm{M})$
 where $\bm{\phi}_{\mathrm{a}}$ can be parameterized as widely adopted SlotAttention (Locatello et al., 2020) and its variants, which is some cross attention with $\bm{S}_{0}$ as queries and $\bm{Z}$ as keys and values.
@@
 Meanwhile, OCL quantizer $\bm{\phi}_{\mathrm{q}}$ transforms $\bm{X}$ into the reconstruction target $\bm{Q}\in\mathbb{R}^{h\times w\times c}$ for the following decoding:
-$\bm{\phi}_{\mathrm{q}}:\bm{X}\rightarrow\bm{Q}$ (3)
+(3) $\bm{\phi}_{\mathrm{q}}:\bm{X}\rightarrow\bm{Q}$
 where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) module (Im Im et al., 2017; Van Den Oord et al., 2017).
@@
 Thirdly, OCL decoder $\bm{\phi}_{\mathrm{d}}$ transforms $\bm{S}$ into reconstruction $\bm{Q}^{\prime}\in\mathbb{R}^{h\times w\times c}$ with destructed $\bm{Q}$ as the condition:
-$\bm{\phi}_{\mathrm{d}}:(\bm{S},\bm{Q})\rightarrow\bm{Q}^{\prime}$ (4)
-where $\bm{\phi}_{\mathrm{d}}$ can be parameterized as a CNN or MLP if using mixture decoding for OCL (Kipf et al., 2022; Seitzer et al., 2023), where each of $\bm{S}$ is spatially broadcast (Watters et al., 2019) into the shape of $\bm{Q}$ and then decoded into a component and finally mixed together with predicted weights; or (ii) a Transformer decoder if using auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where causally masked $\bm{Q}$ is the query and $\bm{S}$ is the key and value; or (iii) a conditional Diffusion model if using diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ with noise is the input and $\bm{S}$ is the condition.
-Reconstructing the destructed $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible.
+(4) $\bm{\phi}_{\mathrm{d}}:(\bm{S},\mathrm{destruct}(\bm{Q}))\rightarrow\bm{Q}^{\prime}$
+Here $\bm{\phi}_{\mathrm{d}}$ can be parameterized as (i) a CNN or MLP for mixture decoding (Kipf et al., 2022; Seitzer et al., 2023), where $\bm{Q}$ is destructed to height and width, and $\bm{S}$ is spatially broadcast (Watters et al., 2019) into this shape and then decoded into components being mixed together; (ii) or a Transformer decoder for auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where $\bm{Q}$ is destructed with causal masking as the query and $\bm{S}$ is the key and value; (iii) or a conditional Diffusion model for diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ is destructed with noise as the input and $\bm{S}$ is the condition.
+Reconstructing $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible.
 Thus, a good reconstruction target, like in (Zhao et al., 2024a, b), is very important.
@@
 Lastly, the supervision signal for OCL training comes from minimizing the reconstruction loss between $\bm{Q}^{\prime}$ and $\bm{Q}$:
-$\mathrm{min}_{(\bm{\phi}_{\mathrm{a}},\bm{\phi}_{\mathrm{d}})}f_{\mathrm{recon}}(\bm{Q}^{\prime},\bm{Q})$ (5)
+(5) $\mathrm{min}_{(\bm{\phi}_{\mathrm{a}},\bm{\phi}_{\mathrm{d}})}f_{\mathrm{recon}}(\bm{Q}^{\prime},\bm{Q})$
 where $f_{\mathrm{recon}}(\cdot,\cdot)$ can be (i) Mean Squared Error (MSE) loss for mixture and diffusion OCL decoding, or (ii) Cross Entropy (CE) loss for auto-regressive decoding.
-3.2 Fully Utilize VFMs in OCL
-We fully utilize VFMs in the OCL setting by two key designs: (i) direct VFM feature extraction to facilitate object-level information aggregation, and (ii) shared VFM feature quantization as reconstruction targets to strengthen supervision in OCL training.
-Figure 3: Better objectness helps OCL aggregation.
-Green and orange areas stand for objects $\bm{o}_{1}$ and $\bm{o}_{2}$, where $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary super-pixels and $\bm{s}_{*}$ is $\bm{o}_{2}$’s centroid.
-But the actual query $\bm{s}\sim\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2})$.
-In VFM super-pixel space, the distance $d_{12}$ between $\bm{v}_{1}$ and $\bm{v}_{2}$ is larger, i.e., better objectness, thus $\bm{s}$ has higher probability $p(d_{s1}>d_{s2})$ to represent $\bm{o}_{2}$ correctly, compared with that in non-VFM super-pixel space.
+3.2. Utilize VFMs in OCL
+In our unified architecture, we utilize VFMs as following.
 Direct VFM Feature Extraction for Better Aggregation
@@
 Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly.
-Thus, We expect the subsequent $\bm{\phi}_{\mathrm{a}}$ to aggregate dense $\bm{Z}$ into sparse slots $\bm{S}$ under queries $\bm{S}_{0}$ better.
-Direct use of VFMs for feature extraction contributes to better aggregation.
-As shown in Figure 1, VFM representations show better objectness or object separability than that of non-VFMs, under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.clust
-er.KMeans.html
-.
-OCL aggregation based on SlotAttention and its variants is essentially a clustering whose initial centroids are trainable (Jia et al., 2023).
-This has already been empirically exploited in previous methods like DINOSAUR (Seitzer et al., 2023) and SlotDiffusion (Wu et al., 2023b), but without any explanation.
-We provide math analysis below.
-Math Analysis: VFM Representation Eases Aggregation
-As shown in Figure 3, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).
-Denote the ideal query of $\bm{o}_{2}$ as $\bm{s}_{*}$, which is the clustering center or centroid of $\bm{o}_{2}$ and is closer to all super-pixels in $\bm{o}_{2}$ than in $\bm{o}_{1}$:
-$d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$ (6)
-where $d(\cdot,\cdot)$ is a distance metric, e.g., minus inner product; $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary points in $\bm{o}_{1}$ and $\bm{o}_{2}$, respectively.
-But the actual query $\bm{s}$ follows $\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2}\bm{I})$. Substituting $\bm{s}$ for $\bm{s}_{*}$, the probability of correct aggregation is:
-[…]
-where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation plane between $\bm{o}_{1}$ and $\bm{o}_{2}$. The closer the boundary is to $\bm{s}_{*}$ the smaller $p_{2}$ is.
-According to Figure 1, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space.
-This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space.
-Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space.
-$\square$
-Figure 4: Qualitative object discovery performance comparison.
+As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html.
+OCL aggregation is essentially a clustering whose initial centroids are trainable (Jia et al., 2023).
+Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$.
+Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason.
+In Section 5.1, we analyze this both theoretically and empirically.
 Shared VFM Feature Quantization for Better Supervision
@@
 In vector quantization, we project $\bm{T}$ with a pre-trainable linear layer and match it with the adjusted $\bm{Z}$:
-$\bm{T}:=\bm{W}\cdot\mathrm{sg}(\bm{T}_{0})$ (8)
-$\bm{D}=||\bm{Z}-\bm{T}||_{2}^{2}$ (9)
+(6) $\bm{T}:=\bm{W}\cdot\mathrm{sg}(\bm{T}_{0})$
+(7) $\bm{D}=||\bm{Z}-\bm{T}||_{2}^{2}$
 where $\mathrm{sg}(\cdot)$ is stop-gradient; $\bm{T}\in\mathbb{R}^{m\times c}$ is the codebook to be used for quantizing $\bm{Z}$; $\bm{D}\in\mathbb{R}^{h\times w\times m}$ is the matching distance between every super-pixel and every code.
 We convert distances to probabilities and select the most matched codes to form the quantization $\bm{Q}$:
-$\bm{P}=\mathrm{softmax}_{c}(-\bm{D})$ (10)
-$\bm{I}=\mathrm{argmax}_{m}(\bm{P})$ (11)
-$\bm{Q}=\mathrm{index}_{m}(\bm{T},\bm{I})$ (12)
+(8) $\bm{P}=\mathrm{softmax}_{c}(-\bm{D})$
+(9) $\bm{I}=\mathrm{argmax}_{m}(\bm{P})$
+(10) $\bm{Q}=\mathrm{index}_{m}(\bm{T},\bm{I})$
 where $\mathrm{softmax}(\cdot)$ is calculated along the channel dimension; $\bm{P}$ is the match probabilities; $\bm{I}\in\mathbb{R}^{h\times w}$ is the matched code indexes; $\mathrm{argmax}(\cdot)$ is calculated along the channel dimension; $\mathrm{index}(\cdot,\cdot)$ is operated along code number dimension.
 The typical STE (Bengio et al., 2013) on $\bm{Q}$, needed in pre-training, can be skipped during OCL training.
-For its pre-training, we introduce some tricks. We add noise to $\bm{D}$ pre-Equation 10 to encourage code utilization:
-$\bm{D}:=\frac{\bm{D}+\bm{G}}{\tau}$ (13)
+For its pre-training, we introduce some tricks. We add noise to $\bm{D}$ before Equation 8 to encourage code utilization:
+(11) $\bm{D}:=\frac{\bm{D}+\bm{G}}{\tau}$
 where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is the Gumbel noise and $\tau$ is the temperature.
-We devise annealing residual connection (Zhao et al., 2024a, b) post-Equation 12 to stabilize pre-training:
-$\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$ (14)
+We devise annealing residual connection (Zhao et al., 2024a, b) after Equation 10 to stabilize pre-training:
+(12) $\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$
 where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing.
 Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$’s elements to be normalized:
-$l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$ (15)
+(13) $l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
 where $\lambda$ is empirically set to 0.1; $\mathbb{E}$ and $\mathbb{V}$ are calculated along height, width and channel dimensions.
@@
 In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) use separate VAE and OCL encoders, which as shown in Figure 1 causes a distribution gap between $\bm{Q}$ and $\bm{Z}$.
-Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen supervision.
-Math Analysis: Shared VFM Representation Quantization Reduces Supervision Noise
-We reconstruct $\bm{Q}^{\prime}$ to approximate the target $\bm{Q}$ ultimately from $\bm{Z}$ via $\bm{\phi}_{\mathrm{a}}\circ\bm{\phi}_{\mathrm{d}}$, denoted as $f$ for simplicity.
-Under MSE loss, the gradient with respect to $\bm{Z}$ is:
-$\frac{\partial\mathrm{MSE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=2(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))\frac{\partial\bm{Q}^{\prime}}{\partial\bm{Z}}$ (16)
-We obtain $\bm{Q}$ by quantizing $\bm{Z}$, i.e., $\mathbb{E}[\bm{Z}]=\bm{Q}$, implying that any deviation of $\bm{Q}^{\prime}=f(\bm{Z})$ from $\bm{Q}$ is due to $f$ and $\bm{Z}$.
-Assuming $f$ preserves $\bm{Z}$’s statistical properties, we have:
-$\mathbb{E}[\bm{Q}^{\prime}]=\mathbb{E}[f(\bm{Z})]\approx\bm{Q}$ (17)
-Thus the residual error $\bm{Q}^{\prime}-\mathrm{sg}({\bm{Q}})$ in Equation 16 is statistically unbiased and small on average:
-$\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q})]\approx 0$ (18)
-But if instead of sharing $\bm{\phi}_{\mathrm{e}}$, we use an extra VAE encoder plus $\bm{\phi}_{\mathrm{a}}$ to obtain the target, denoted as $\bm{Q}_{2}$, then $\bm{Q}_{2}\neq\bm{Q}$.
-Substitute $\bm{Q}_{2}$ into Equation 16 and the residual error $\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})$ would be systematically biased:
-$\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})]\neq 0$ (19)
-which increases noise in optimization.
-Using CE loss is slightly different but it does not alter our judgment above. $\square$
-resolution= ClevrTex #slot=11 COCO (panoptic) #slot=7 VOC #slot=6 MOVi-D #slot=21, conditional
-256$\times$256 ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO
-SLATE-DINO 15.8
-$\pm$0.3
-85.8
-$\pm$0.3
+Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision.
+In Section 5.2, we analyze this both theoretically and empirically.
+3.3. Compare Architectures
+Figure 3. Model architecture comparison.
+As shown in Figure 3, we compare the baselines with our VVO in a unified perspective. Specifically,
+- Our VVO: (1) VFMs are employed for OCL encoding and their features are fed to the aggregator directly, which eases OCL aggregation, as formulated in Section 3.2.
+(2) VFM features are shared quantized as reconstruction targets, which strengthens OCL supervision, as formulated in Section 3.2.
+- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator.
+(2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter.
+This loses much information of VFM features, while shared quantization can preserve such information (Van Den Oord et al., 2017).
+- SLATE (the unofficially improved version; we adopt it) (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b) / STEVE (Singh et al., 2022c): (1) Same as our VVO.
+(2) Reconstruction targets are discretized from features of a separate VAE encoder, instead of quantized from features shared from the same VFM for OCL encoding, which causes optimization noises.
+- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as our VVO.
+(2) Reconstruction targets are quantized from features of a separate VAE encoder, instead of sharing the same VFM for OCL encoding, which causes optimization noises.
+- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as our VVO.
+(2) Reconstruction targets are shared from the same VFM for OCL encoding but without quantization, which causes optimization noises.
+Our VVO adopts a clean and unified architecture with the above-mentioned two key designs.
+The first design, direct VFM feature extraction for OCL aggregation, is not proposed by us but we are the first to provide both theoretical and empirical analyses.
+The second design, shared VFM feature quantization as reconstruction targets, is proposed by us.
+Such shared quantization is also the key point that our VVO supports various OCL decoders.
+Specifically,
+- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features.
+As quantized features are actually the clustering centroids of continuous features, these decoders are inherently compatible with our shared quantized VFM features.
+- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are originally designed for discretized features, while (Seitzer et al., 2023) and (Kakogeorgiou et al., 2024) also explore their applicability to continuous features.
+Thus these decoders are naturally applicable to our shared quantized VFM features.
+- Diffusion-based OCl decoders, e.g., Conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), are originally designed for dicretized features.
+Thus if without our shared quantized VFM features as reconstruction targets, then this type of OCL decoders would not be unified into our architecture.
+Figure 4. Qualitative object discovery performance comparison.
+resolution= ClevrTex #slot=11 COCO #slot=7 VOC #slot=6 MOVi-D #slot=21 conditional
+256$\times$256 (224) ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
+SLATE-DINO 17.4
+±2.9
+87.4
+±1.7
+44.5
+±2.2
+43.3
+±2.4
+17.5
+±0.6
+28.8
+±0.3
+26.8
+±0.3
+25.4
+±0.3
+18.6
+±0.1
 26.2
-$\pm$0.1
-40.0
-$\pm$0.2
-34.4
-$\pm$0.3
-35.4
-$\pm$0.5
-7.1
-$\pm$0.1
-26.8
-$\pm$0.7
-16.4
-$\pm$0.6
-22.2
-$\pm$0.9
-13.0
-$\pm$0.4
-34.7
-$\pm$0.9
+±0.8
+37.2
+±0.5
+36.1
+±0.4
 VQDINO
 Tfd
-21.5 $\pm$0.2 86.1 $\pm$0.1 29.0 $\pm$0.1 43.3 $\pm$0.1 38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4 19.0 $\pm$0.9 24.1 $\pm$0.7 14.3 $\pm$0.3 37.2 $\pm$0.8
-STEVE-DINO 35.0
-$\pm$0.9
-64.1
-$\pm$2.9
-21.2
-$\pm$0.7
-23.2
-$\pm$1.2
+55.4 ±18.2 85.9
+±0.7
+54.4 ±2.4 53.6 ±2.5 21.1 ±2.1 31.5 ±1.1 29.6 ±0.7 28.2 ±0.8 22.7 ±0.9 26.6 ±1.3 39.8 ±0.3 38.9 ±0.3
+STEVE-DINO 33.7
+±0.2
+66.2
+±0.3
+22.6
+±0.2
+20.9
+±0.3
 VQDINO
 TfdT
-41.4 $\pm$0.8 70.5 $\pm$3.3 23.1 $\pm$0.6 25.9 $\pm$1.4
-DINOSAUR-DINO 18.5
-$\pm$0.2
-87.8
-$\pm$0.4
-26.7
-$\pm$0.3
-39.9
-$\pm$0.5
-35.3
-$\pm$0.1
-36.4
-$\pm$0.2
-7.1
-$\pm$0.0
-27.3
-$\pm$0.2
-18.8
-$\pm$0.4
-31.2
-$\pm$0.1
-14.4
-$\pm$0.2
-37.9
-$\pm$0.4
+37.2 ±0.7 72.7 ±4.9 25.7 ±1.5 24.3 ±1.7
+DINOSAUR-DINO 50.7
+±24.1
+89.4
+±0.3
+53.3
+±5.0
+52.8
+±5.2
+18.2
+±1.0
+37.0
+±1.2
+28.3
+±0.5
+26.9
+±0.5
+21.5
+±0.7
+36.2
+±1.3
+40.6
+±0.6
+39.7
+±0.6
 VQDINO
 Mlp
-51.2 $\pm$0.4 88.3 $\pm$0.3 29.8 $\pm$0.2 44.6 $\pm$0.4 35.4 $\pm$0.6 36.7 $\pm$0.5 7.6 $\pm$0.1 27.7 $\pm$0.1 20.6 $\pm$0.1 32.9 $\pm$0.1 15.6 $\pm$0.0 39.1 $\pm$0.3
-SlotDiffusion-DINO 60.1
-$\pm$9.4
-81.6
-$\pm$2.7
-31.4
-$\pm$0.9
-45.3
-$\pm$0.9
-34.9
-$\pm$0.4
-36.0
-$\pm$0.4
-7.0
-$\pm$0.2
-26.7
-$\pm$0.6
-15.0
-$\pm$0.3
-19.7
-$\pm$0.4
-12.3
-$\pm$0.2
-33.1
-$\pm$0.3
+66.0 ±0.3 88.7
+±0.6
+57.0 ±0.5 56.6 ±0.5 19.2 ±0.4 35.9
+±0.6
+28.7 ±0.3 27.4 ±0.3 21.7 ±0.8 35.4
+±1.0
+41.0 ±0.1 40.1 ±0.2
+SlotDiffusion-DINO 66.1
+±1.3
+82.7
+±1.6
+54.3
+±0.5
+53.4
+±0.8
+17.7
+±0.5
+29.0
+±0.1
+27.0
+±0.4
+25.6
+±0.4
+17.0
+±1.2
+21.7
+±1.8
+35.2
+±0.9
+34.0
+±1.0
 VQDINO
 Dfz
-37.9 $\pm$2.1 81.7 $\pm$2.6 29.3 $\pm$1.1 47.8 $\pm$1.4 37.2 $\pm$0.7 38.3 $\pm$0.9 7.3 $\pm$0.1 27.8 $\pm$0.9 17.8 $\pm$0.1 24.1 $\pm$0.3 14.1 $\pm$0.1 36.1 $\pm$0.3
-Table 1: Object discovery performance with DINO2 ViT (s/14) for OCL encoding.
+72.2 ±0.2 81.9
+±2.0
+57.6 ±0.7 56.8 ±0.8 18.3 ±0.4 28.7
+±1.0
+27.2 ±0.1 25.8 ±0.1 19.1 ±0.5 24.2 ±0.3 36.9 ±0.1 35.9 ±0.2
+Table 1.
+Object discovery performance with DINO2 ViT (s/14) for OCL encoding.
 VVO is instantiated as VQDINO; Tfd, TfdT, Mlp and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
-resolution= ClevrTex #slot=11 COCO (panoptic) #slot=7 VOC #slot=6 MOVi-D #slot=21, conditional
-256$\times$256 ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO
-SLATE-SAM 12.4
-$\pm$1.1
-80.3
-$\pm$0.6
-28.4
-$\pm$0.3
-44.0
-$\pm$0.8
-34.7
-$\pm$0.6
-35.8
-$\pm$1.4
-6.4
-$\pm$0.3
-25.9
-$\pm$0.4
-12.7
-$\pm$0.4
-19.5
-$\pm$0.9
-11.7
-$\pm$0.2
-31.6
-$\pm$1.1
+resolution= ClevrTex #slot=11 COCO #slots=7 VOC #slot=6 MOVi-D #slot=21 conditional
+256$\times$256 ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
+SLATE-SAM 18.0
+±0.3
+82.8
+±0.9
+51.8
+±0.8
+50.6
+±0.8
+12.5
+±0.8
+27.4
+±1.4
+22.7
+±0.8
+21.5
+±0.8
+16.3
+±0.7
+21.4
+±1.3
+34.8
+±0.5
+33.9
+±0.6
 VQSAM
 Tfd
-14.2 $\pm$1.6 81.1 $\pm$0.2 28.5 $\pm$0.4 44.8 $\pm$0.6 36.5 $\pm$0.8 37.9 $\pm$0.8 7.4 $\pm$0.1 27.9 $\pm$0.5 17.3 $\pm$0.6 19.6 $\pm$0.7 13.1 $\pm$0.2 34.5 $\pm$0.8
-STEVE-SAM 28.0
-$\pm$0.5
-59.3
-$\pm$1.2
-18.0
-$\pm$0.3
-20.3
-$\pm$0.9
+19.9 ±4.1 87.3 ±1.1 53.3 ±2.4 51.6 ±2.8 16.4 ±0.7 26.3
+±0.5
+26.6 ±0.5 25.2 ±0.6 18.6 ±1.0 22.2 ±1.8 37.2 ±0.6 36.0 ±0.6
+STEVE-SAM 31.2
+±3.8
+62.9
+±5.1
+22.2
+±2.2
+19.9
+±2.4
 VQSAM
-Ttdt
-39.7 $\pm$0.2 68.5 $\pm$0.8 24.2 $\pm$0.2 26.3 $\pm$0.4
-DINOSAUR-SAM 14.4
-$\pm$0.3
-87.6
-$\pm$0.9
-25.8
-$\pm$0.7
-39.3
-$\pm$0.7
-27.3
-$\pm$0.4
-27.7
-$\pm$0.5
-5.9
-$\pm$0.1
-22.6
-$\pm$0.5
-11.8
-$\pm$0.7
-16.7
-$\pm$1.1
-11.2
-$\pm$0.3
-29.6
-$\pm$1.2
+TfdT
+31.3 ±3.7 63.8 ±2.5 22.6 ±1.3 20.5 ±1.4
+DINOSAUR-SAM 21.1
+±0.8
+88.7
+±0.6
+46.3
+±0.5
+45.3
+±0.5
+11.7
+±2.6
+21.6
+±1.4
+21.8
+±2.5
+20.6
+±2.5
+12.1
+±0.4
+15.6
+±1.8
+29.4
+±0.6
+28.7
+±0.6
 VQSAM
 Mlp
-25.3 $\pm$2.6 84.1 $\pm$0.2 29.4 $\pm$1.1 44.1 $\pm$0.9 36.5 $\pm$0.8 37.6 $\pm$0.8 7.2 $\pm$0.1 27.6 $\pm$0.3 17.3 $\pm$0.6 19.2 $\pm$0.7 12.8 $\pm$0.4 32.2 $\pm$0.8
-SlotDiffusion-SAM 19.0
-$\pm$1.0
-76.7
-$\pm$2.3
-31.2
-$\pm$0.5
-47.3
-$\pm$1.2
-34.5
-$\pm$0.3
-35.6
-$\pm$0.3
-7.0
-$\pm$0.1
+64.1 ±1.1 87.4
+±0.7
+58.8 ±0.7 58.0 ±0.8 13.2 ±0.3 25.2 ±1.2 23.5 ±0.4 22.2 ±0.3 14.2 ±0.1 19.4 ±0.6 32.6 ±0.2 31.6 ±0.3
+SlotDiffusion-SAM 19.5
+±0.5
+81.2
+±0.6
+55.1
+±0.4
+53.6
+±0.4
+16.8
+±0.5
+27.1
+±0.7
 26.7
-$\pm$0.3
-15.2
-$\pm$0.6
-17.2
-$\pm$1.3
-12.6
-$\pm$0.7
-33.9
-$\pm$1.6
+±0.4
+25.3
+±0.4
+17.0
+±0.9
+18.9
+±2.1
+36.2
+±0.6
+34.9
+±0.8
 VQSAM
 Dfz
-16.7 $\pm$0.3 76.9 $\pm$0.3 32.5 $\pm$0.7 49.2 $\pm$0.9 35.7 $\pm$0.6 36.8 $\pm$0.6 7.2 $\pm$0.0 27.1 $\pm$0.4 16.9 $\pm$1.3 19.6 $\pm$1.4 13.6 $\pm$0.2 36.0 $\pm$0.9
-Table 2: Object discovery performance with SAM2 Hiera+FPN (t/16) for OCL decoding.
+30.0 ±0.3 78.6
+±0.8
+57.3 ±0.2 56.1 ±0.2 17.3 ±0.4 26.0
+±0.2
+27.2 ±0.4 25.8 ±0.4 17.3 ±0.5 20.2 ±2.3 36.3 ±0.8 35.1 ±0.9
+Table 2.
+Object discovery performance with SAM2 Hiera+FPN (t/16) for OCL decoding.
 VVO is instantiated as VQSAM; Tfd, TfdT, Mlp and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
-4 Experiment
+4. Experiment
 We conduct all experiments using three random seeds.
-4.1 Set up the Benchmark
+4.1. Set up the Benchmark
 Datasets.
@@
 ClevrTex [^2] https://www.robots.ox.ac.uk/~vgg/data/clevrtex comprises synthetic images, each with about 10 geometric objects scattered in complex background.
-MOVi-D [^3] https://github.com/google-research/kubric/blob/main/challeng
-es/movi/README.md#movi-d
+MOVi-D [^3] https://github.com/google-research/kubric/blob/main/challenges/movi/README.md#
+movi-d
 contains synthetic videos, each with up to 20 daily objects dropping and bumping.
@@
 VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation.
-We also report results on real-world video dataset YTVIS [^6] https://youtube-vos.org/dataset/vis version HQ [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube.
+We also report results on real-world video dataset YTVIS version HQ [^6] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube.
 We choose Physion [^7] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
@@
 Instead of copying existing results, we reproduce all baselines to realize fair comparison.
-As detailed in Appendices A and B, we use identical data augmentation, VFMs in OCL encoding and training recipes for all experiment items unless not applied.
-We instantiate all baselines’ VAE part as TAESD [^9] https://huggingface.co/docs/diffusers/en/api/models/autoenco
-der_tiny
-, which is a large-scale pretrained StableDiffusion [^10] https://huggingface.co/spaces/stabilityai/stable-diffusion module, to build all strong baselines.
-ARI ARIfg mIoU mBO
-Using higher resolution: 384$\times$384
-resolution=384$\times$384 COCO (panoptic) #slot=7
-SLATE-DINO 50.9
-$\pm$0.3
-51.0
-$\pm$0.9
-9.0
-$\pm$0.1
-27.9
-$\pm$0.7
+We use identical data augmentation, VFMs in OCL encoding and training recipes for all experiment items unless not applied.
+We instantiate all baselines’ VAE part as TAESD [^8] https://huggingface.co/docs/diffusers/en/api/models/autoencoder_tiny, which is a large-scale pretrained StableDiffusion [^9] https://huggingface.co/spaces/stabilityai/stable-diffusion module, to build all strong baselines.
+ARI ARIfg mBO mIoU
+Using higher resolution: 384$\times$384 (336)
+resolution=384$\times$384 (336) COCO #slot=7
+SLATE-DINO 41.4
+±1.0
+34.0
+±0.3
+27.4
+±0.4
+25.9
+±0.5
 VQDINO
 Tfd
-55.6 $\pm$0.4 55.9 $\pm$0.3 9.8 $\pm$0.0 30.2 $\pm$0.2
-DINOSAUR-DINO 53.8
-$\pm$0.2
-54.1
-$\pm$0.4
-9.2
-$\pm$0.0
-28.9
-$\pm$0.3
+44.1 ±0.8 37.5 ±1.1 29.6 ±0.5 28.0 ±0.5
+DINOSAUR-DINO 45.0
+±0.1
+42.2
+±0.5
+29.9
+±0.1
+28.5
+±0.1
 VQDINO
 Mlp
-54.7 $\pm$0.3 55.2 $\pm$0.6 9.4 $\pm$0.1 29.8 $\pm$0.3
-SlotDiffusion-DINO 53.1
-$\pm$0.6
-53.3
-$\pm$0.8
-9.1
-$\pm$0.2
-28.4
-$\pm$0.7
+44.6
+±0.7
+42.6 ±0.5 29.8
+±0.3
+28.6 ±0.3
+SlotDiffusion-DINO 41.6
+±0.5
+34.5
+±0.4
+27.7
+±0.2
+26.2
+±0.2
 VQDINO
 Dfz
-55.3 $\pm$0.8 55.7 $\pm$0.5 9.8 $\pm$0.1 30.2 $\pm$0.4
+43.4 ±1.3 34.2
+±0.4
+28.3 ±0.7 26.9 ±0.7
 Using different aggregators: SlotAttention, BO-QSA
-resolution=256$\times$256 COCO (panoptic) #slot=7
-SLATE-DINO-SlotAttention 34.1
-$\pm$0.7
-34.9
-$\pm$0.6
-7.0
-$\pm$0.1
-26.2
-$\pm$0.9
-VQDINO Tfd -SlotAttention 37.7 $\pm$0.5 39.3 $\pm$0.4 7.7 $\pm$0.1 28.0 $\pm$0.6
-SLATE-DINO-BO-QSA 34.4
-$\pm$0.3
-35.4
-$\pm$0.5
-7.1
-$\pm$0.1
+resolution=256$\times$256 (224) COCO #slot=7
+SLATE-DINO-SlotAttention 17.0
+±1.3
+28.3
+±0.5
+26.4
+±0.4
+25.1
+±0.3
+VQDINO Tfd -SlotAttention 20.8 ±2.0 31.5 ±1.2 29.4 ±0.9 27.9 ±1.1
+SLATE-DINO-BO-QSA 17.5
+±0.6
+28.8
+±0.3
 26.8
-$\pm$0.7
-VQDINO Tfd -BO-QSA 38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4
-Table 3: VVO using higher resolution (upper) and different aggregators (lower) on object discovery. By default, we use BO-QSA for all our experiment items, including the baselines.
-ARI ARIfg mIoU mBO
+±0.3
+25.4
+±0.3
+VQDINO Tfd -BO-QSA 21.1 ±2.1 31.5 ±1.1 29.6 ±0.7 28.2 ±0.8
+Table 3. VVO using higher resolution (upper) and different aggregators (lower) on object discovery. By default, we use BO-QSA for all our experiment items, including the baselines.
+ARI ARIfg mBO mIoU
 Compared with general improvers: GDR and MSF
-resolution=256$\times$256 COCO (panoptic) #slot=7
+resolution=256$\times$256 (224) COCO #slot=7
 GDR
@@
 -DINO
-34.8
-$\pm$0.4
-36.0
-$\pm$0.6
-7.1
-$\pm$0.1
-27.1
-$\pm$0.7
+18.0
+±1.4
+29.2
+±0.2
+27.4
+±0.7
+26.0
+±0.7
 MSF
@@
 -DINO
-35.9
-$\pm$0.3
-37.1
-$\pm$0.4
-7.2
-$\pm$0.0
+18.0
+±0.5
+29.0
+±0.2
 27.4
-$\pm$0.5
+±0.3
+26.1
+±0.3
 VQDINO
 Tfd
-38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4
+21.1 ±2.1 31.5 ±1.1 29.6 ±0.7 28.2 ±0.8
 GDR
@@
 -DINO
-36.2
-$\pm$0.5
-37.1
-$\pm$0.4
-7.1
-$\pm$0.1
+17.9
+±0.1
+29.0
+±0.3
 27.2
-$\pm$0.6
+±0.1
+25.8
+±0.1
 MSF
@@
 -DINO
-35.6
-$\pm$0.2
-36.6
-$\pm$0.4
-7.1
-$\pm$0.0
-27.2
-$\pm$0.4
+16.9
+±0.4
+28.7
+±0.1
+26.6
+±0.2
+25.2
+±0.2
 VQDINO
 Dfz
-37.2 $\pm$0.7 38.3 $\pm$0.9 7.3 $\pm$0.1 27.8 $\pm$0.9
+18.3 ±0.4 28.7 ±1.0 27.2 ±0.1 25.8 ±0.1
 Compared with SotA methods: SPOT, VideoSAUR
-resolution=256$\times$256 COCO (panoptic) #slot=7
-SPOT-DINO 37.5
-$\pm$0.4
-38.7
-$\pm$0.3
-7.8
-$\pm$0.1
-28.1
-$\pm$0.9
+resolution=256$\times$256 (224) COCO #slot=7
+SPOT-DINO 20.3
+±0.7
+41.1
+±0.3
+30.4
+±0.1
+29.0
+±0.9
 VQDINO
 Tfd9
-+self-training
-39.3 $\pm$0.3 40.4 $\pm$0.4 8.0 $\pm$0.0 29.6 $\pm$0.5
-resolution=256$\times$256 YTVIS (HQ) #slot=7, unconditional
-VideoSAUR-DINO 34.0
-$\pm$0.6
-39.1
-$\pm$0.9
-27.2
-$\pm$0.4
-32.2
-$\pm$0.6
+21.3 ±0.4 42.3 ±1.0 31.4 ±0.2 29.9 ±0.3
+resolution=256$\times$256 (224) YTVIS (HQ) #slot=7, unconditional
+VideoSAUR-DINO 33.0
+±0.6
+49.0
+±0.9
+30.8
+±0.4
+30.1
+±0.6
 VQDINO
-Smd
-40.7 $\pm$0.5 43.5 $\pm$0.6 29.2 $\pm$0.2 35.1 $\pm$0.5
-Table 4: VVO compared with general improvers (upper) and SotA methods (lower) on object discovery.
-SPOT uses both Transformer decoder with 9 permutations Tfd9 and “self-training”; VideoSAUR uses SlotMixer Smd as its decoder.
-4.2 Evaluate on Object Discovery
+SmdT
+35.7 ±0.5 49.5 ±0.6 32.7 ±0.2 31.6 ±0.5
+Table 4. VVO versus general improvers (upper) and SotA methods (lower) on object discovery. SPOT uses Transformer with 9 permutations Tfd9 as decoder while VideoSAUR uses SlotMixer SmdT.
+4.2. Evaluate on Object Discovery
 Object discovery task intuitively shows how well those slots separate different objects.
-We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metri
-cs.adjusted_rand_score.html
-(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metri
-cs.jaccard_score.html
-(mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
+We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^10] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_sco
+re.html
+(ARI), ARIfg (foreground), mean Intersection-over-Union [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
 With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO.
@@
 VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
-4.3 Evaluate on Set Prediction
+4.3. Evaluate on Set Prediction
 class labels bounding boxes
@@
 DINOSAUR + MLP 0.61
-$\pm$0.0
+±0.0
 0.57
-$\pm$0.1
+±0.1
 VQDINO
@@
 + MLP
-0.64 $\pm$0.1 0.59 $\pm$0.1
-Table 5: Set prediction performance on COCO (panoptic, #slot=7).
+0.64 ±0.1 0.59 ±0.1
+Table 5. Set prediction performance on COCO (panoptic, #slot=7).
 Set prediction task directly shows how much object information those slots grasp.
 We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023).
-We measure top1 accuracy of the classified class labels while we measure the R2 score [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metri
-cs.r2_score.html
-of the regressed bounding box coordinates.
-As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO-MLP, obtains better object classification and bounding box regression.
+We measure top1 accuracy of the classified class labels while we measure the R2 score [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.r2_score.html of the regressed bounding box coordinates.
+As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO Mlp , obtains better object classification and bounding box regression.
 Thus, our models’ object representations are of higher quality.
-4.4 Deploy to the Downstream
-Figure 5: Visual prediction (upper) and reasoning (lower) performance on Physion (#slot=8). VVO has smaller prediction error in all time steps, and higher reasoning accuracy in later time steps.
+4.4. Deploy to the Downstream
+Figure 5. Visual prediction (upper) and reasoning (lower) performance on Physion (#slot=8). VVO has smaller prediction error in all time steps, and higher reasoning accuracy in later time steps.
 Better object representation benefits downstream tasks.
@@
 Then the object-centric dynamics model SlotFormer (Wu et al., 2023a) is trained on those slots in an auto-regressive manner along the time dimension.
-We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO-MLP to extract slots.
+We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO Mlp to extract slots.
 On visual prediction, we evaluate the prediction errors per time step measured in normalized MSE between regressed and extracted slots.
@@
 As shown in Figure 5 lower, VVO accuracies are slightly lower at the beginning but much higher later than the baseline.
-4.5 Ablate the Architecture
+4.5. Ablate the Architecture
 As shown in Table 6, VVO’s design of shared VAE and OCL encoder consistently outperforms separate VAE and OCL encoders, even when the latter employs another VFM for VAE encoding.
@@
 ARI ARIfg
-shared DINO2 ViT DINO2 ViT 38.5
-$\pm$0.4
-39.6
-$\pm$0.4
-separate TAESD encoder ResNet18 28.3 $\pm$1.3 29.1 $\pm$1.5
-TAESD encoder DINO2 ViT 34.6
-$\pm$0.7
-35.8
-$\pm$0.5
-SAM2 encoder DINO2 ViT 34.9
-$\pm$0.3
-35.8
-$\pm$0.4
-Table 6: VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders.
+shared DINO2 ViT DINO2 ViT 21.1
+±2.1
+31.5
+±1.1
+separate TAESD encoder ResNet18 15.4 ±3.3 24.1 ±2.5
+TAESD encoder DINO2 ViT 17.5
+±0.6
+28.8
+±0.3
+SAM2 encoder DINO2 ViT 18.3
+±1.3
+29.1
+±0.9
+Table 6.
+VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders.
 Results are on COCO panoptic.
@@
 OCL encoder noise residual regulariz. ARI ARIfg
-37.2
-$\pm$0.7
-38.3
-$\pm$0.9
-35.4
-$\pm$2.5
-36.6
-$\pm$2.2
-36.5
-$\pm$1.6
-37.6
-$\pm$1.9
-36.8
-$\pm$1.3
-37.8
-$\pm$1.0
-36.9
-$\pm$0.7
-37.9
-$\pm$0.8
-Table 7: VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts.
+21.1
+±2.1
+31.5
+±1.1
+18.3
+±1.3
+29.1
+±0.9
+19.5
+±0.6
+29.9
+±1.9
+20.1
+±1.3
+29.8
+±1.0
+20.0
+±0.7
+29.4
+±0.8
+Table 7.
+VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts.
 In comparison to VVO’s key designs, these tricks are more like the cherry on top.
 Results are on COCO panoptic with settings consistent with the above.
-Figure 6: Effects of tricks in our VQ variant: (upper) Gumbel noise improves codebook utilization, where “CV” means Coefficient of Variation, and curves are smoothed by Gaussian kernel of size 10; (lower) Annealing residual connection improves VAE pretraining, and the blue curve’s turning point at epoch 4 is where the residual connection anneals to zero.
-Results are from the VAE pretraining of VQDINO-Dfz on COCO panoptic.
-Conclusion
+Figure 6.
+Effects of tricks in our VQ variant: (upper) Gumbel noise improves codebook utilization, where “CV” means Coefficient of Variation, and curves are smoothed by Gaussian kernel of size 10; (lower) Annealing residual connection improves VAE pretraining, and the blue curve’s turning point at epoch 4 is where the residual connection anneals to zero.
+Results are from the VAE pretraining of VQDINO
+Dfz
+on COCO panoptic.
+5. Analysis
+We provide both mathematical modeling and statistical verification on why and how our VVO is better.
+5.1. Aggregation as Clustering
+Figure 7.
+Better objectness helps OCL aggregation.
+Green and orange areas stand for objects $\bm{o}_{1}$ and $\bm{o}_{2}$, where $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary super-pixels and $\bm{s}_{*}$ is $\bm{o}_{2}$’s centroid.
+But the actual query $\bm{s}\sim\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2})$.
+In VFM super-pixel space, the distance $d_{12}$ between $\bm{v}_{1}$ and $\bm{v}_{2}$ is larger, i.e., better objectness, thus $\bm{s}$ has higher probability $p(d_{s1}>d_{s2})$ to represent $\bm{o}_{2}$ correctly, compared with that in non-VFM super-pixel space.
+Mathematic Modeling
+As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).
+Denote the ideal query of $\bm{o}_{2}$ as $\bm{s}_{*}$, which is the clustering center or centroid of $\bm{o}_{2}$ and is closer to all super-pixels in $\bm{o}_{2}$ than in $\bm{o}_{1}$:
+(14) $d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$
+where $d(\cdot,\cdot)$ is a distance metric, e.g., minus inner product; $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary points in $\bm{o}_{1}$ and $\bm{o}_{2}$, respectively.
+But the actual query $\bm{s}$ follows $\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2}\bm{I})$. Substituting $\bm{s}$ for $\bm{s}_{*}$, the probability of correct aggregation is:
+[…]
+where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation plane between $\bm{o}_{1}$ and $\bm{o}_{2}$. The closer the boundary is to $\bm{s}_{*}$ the smaller $p_{2}$ is.
+According to Figure 1 the first observation, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space.
+This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space.
+Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space.
+$\square$
+Statistical Verification
+Equation 15 is very difficult to solve, but we can use Monte Carlo estimation:
+(16) $p_{2}=\frac{1}{N}\Sigma_{i=1}^{N}\mathbbm{1}[\bm{v}_{i}\in\bm{o}_{2}]$
+Specifically, we respectively use ResNet18, DINO2 ViT and SAM2 encoder to extract images into super-pixels, then apply KMeans clustering to find each clusters mean and std.
+From the Gaussian distributions defined by these mean and std, we repeatedly draw samples and calculate the ratio that such samples fall within their original clusters.
+This ratio is the correct aggregation probability.
+Figure 8. Statistical correlation between correct aggregation probability and object discovery accuracy on VOC.
+A step further, we can calculate the statistical correlation between the correct aggregation probability and the corresponding OCL performance.
+As shown in Figure 8, the correlation is roughly significant.
+Considering that correct aggregation probability can be calculated before the expensive OCL training, our Equation 15 is instructive to training better OCL models by selecting OCL encoders.
+5.2. Shared Quantization and Optimization Noise
+Mathematic Modeling
+We reconstruct $\bm{Q}^{\prime}$ to approximate the target $\bm{Q}$ ultimately from $\bm{Z}$ via $\bm{\phi}_{\mathrm{a}}\circ\bm{\phi}_{\mathrm{d}}$, denoted as $f$ for simplicity.
+Under MSE loss, the gradient with respect to $\bm{Z}$ is:
+(17) $\frac{\partial\mathrm{MSE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=2(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))\frac{\partial\bm{Q}^{\prime}}{\partial\bm{Z}}$
+We obtain $\bm{Q}$ by quantizing $\bm{Z}$, i.e., $\mathbb{E}[\bm{Z}]=\bm{Q}$, implying that any deviation of $\bm{Q}^{\prime}=f(\bm{Z})$ from $\bm{Q}$ is due to $f$ and $\bm{Z}$.
+Assuming $f$ preserves $\bm{Z}$’s statistical properties, we have:
+(18) $\mathbb{E}[\bm{Q}^{\prime}]=\mathbb{E}[f(\bm{Z})]\approx\bm{Q}$
+Thus the residual error $\bm{Q}^{\prime}-\mathrm{sg}({\bm{Q}})$ in Equation 17 is statistically unbiased and small on average:
+(19) $\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q})]\approx 0$
+But if instead of sharing $\bm{\phi}_{\mathrm{e}}$, we use an extra VAE encoder plus $\bm{\phi}_{\mathrm{q}}$ to obtain the target, denoted as $\bm{Q}_{2}$, then $\bm{Q}_{2}\neq\bm{Q}$ according to Figure 1 the second observation.
+Substitute $\bm{Q}_{2}$ into Equation 17 and the residual error $\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})$ would be systematically biased:
+(20) $\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})]\neq 0$
+which increases noise in optimization.
+Under CE loss, the gradient with respective to $\bm{Z}$ is:
+(21) $\frac{\partial\mathrm{CE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=\frac{\partial f(\bm{Z})^{T}}{\partial\bm{Z}}(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))$
+where $\frac{\partial f(\bm{Z})}{\partial\bm{Z}}$ is the Jacobian matrix of $f(\bm{Z})$ with respect to $\bm{Z}$.
+Anyway, this has similar structure to Equation 17 and thus does not alter our judgment above.
+$\square$
+Statistical Verification
+Figure 9. Gradient distributions tend to disperse or have outliers if not using our shared VFM feature quantization. Results here are collected at time step 10k, on dataset VOC.
+The statistical bias in Equation 20 manifests as the gradient distribution of different layers, especially that of slots.
+No statistical bias is usually associated with a smooth zero-mean small-std Gaussian gradient distribution.
+Our technique, shared VFM feature quantization as reconstruction targets, includes two key points, sharing and quantization.
+Thus we compare three cases: (i) Shared and quantized, where DINO features are both used for aggregation and quantized for reconstruction; (ii) Shared but not quantized, where DINO features are used for aggregation and direct for reconstruction without quantization; (iii) Not shared but quantized, where DINO features are used for aggregation while quantized SAM features are used for reconstruction.
+As shown in Figure 9, using CE or MSE as the loss, our shared quantization achieves smoother Gaussian gradient distributions, which benefits OCL training.
+If not shared but quantized, then gradients tend to disperse; while if shared but not quantized, then gradients have some outliers.
+6. Conclusion
 We propose a unified architecture VVO for object-centric representation learning.
@@
 VVO has the potential to serve as a general testbed for research related to OCL in the future.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Dataset Processing
-The datasets for image OCL are ClevrTex, COCO (panoptic) and VOC. The datasets for video OCL are MOVi-D and YTVIS. For different datasets, we have mostly the same processing but a bit differences.
-A.1 Shared Preprocessing
-To accelerate experiments, we adopt the dataset conversion scheme: Convert all datasets into LMDB database format, and load it into NVMe disk or RAM disk to minimize I/O overhead and maximize throughput.
-Specifically, all images or video frames are resized by short side to 256, as well as their segmentation masks, and then are stored data type uint8 to save space.
-For images or videos, we use default interpolation mode; but for segmentation masks, we use NEAREST-EXACT interpolation mode.
-During training, for both images and videos, we use spatial resolution of 256$\times$256 as inputs and outputs.
-We normalize the input images or videos by ImageNet-Mean and -Std, which are the common normalization scheme using by VFMs like DINO and SAM.
-For videos, we also using random strided temporal cropping of fixed window size 6 to accelerate training and improve generalization.
-During testing, there is no difference to that of training for images. But for videos, we need to remove the random strided temporal cropping and use the full video of 24 time steps.
-A.2 Dataset-Specific Processing
-ClevrTex. As there are 10 objects at most in each image, we use number of slots queries of 10 + 1, where 1 represents for the background.
-COCO & VOC.
-We use the panoptic segmentation annotations of COCO for evaluation, which is more challenging than instance segmentation.
-Besides, OCL’s queries-based workflow intuitively fits panoptic segmentation rather than instance segmentation or semantic segmentation.
-There can be lots of or only two objects in an image, but we just follow the popular hyperparameter selection, i.e., 7 slots for COCO and 6 slots for VOC, respectively.
-MOVi-D.
-There are abundant annotations like depth, optical-flow, camera intrinsic/extrinsic parameters and so on.
-But for simple comparison, we only utilize the segmentation masks for evaluation and the bounding boxes as slots initialization.
-As there are 23 objects in these dataset, we use number of slots queries of 23+1.
-YTVIS.
-We use the original YTVIS videos and HQ-YTVIS segmentation annotations.
-The latter provides high-quality masks.
-During VAE pretraining, we utilize all the video frames.
-But during OCL training, we only utilize the every-5-time-step annotations for calculating the performance.
-Appendix B Training Recipe
-Following the convention, there are two training stages.
-The pretraining stage is for training VAE modules on corresponding datasets, so as to learn good discrete intermediate representation.
-The OCL training stage, which utilize the pretraining VAE representation as guidance for object centric learning.
-B.1 Pretrain VAE
-For all datasets, we use total number of training iterations 30,000, and validation interval iterations 600, so that we have about 50 checkpoints for every OCL model on every dataset.
-But to save storage, we save only the latter half 25 checkpoints.
-This is unified across all datasets.
-The batch size for image datasets is 64, while for video datasets it is 16. This holds for both training and validation. This setting is shared across all datasets.
-For multi-process, we use the number of workers 4. This holds for both training and validation. This setting is shared across all datasets.
-We use Adam optimizer with initial learning rate of 2e-3.
-And the learning rate is manipulated by cosine annealing scheduling, with a linear warmup during first 1/20 of total steps.
-This setting is shared across all datasets.
-We use automatic mixed precision provided by PyTorch autocast API plus compile.
-Along with this, we use PyTorch builtin gradient scaler, so as to apply gradient clipping of maximum norm 1.0.
-This setting is shared across all datasets.
-B.2 Train OCL Model
-For all datasets, we use total number of training iterations 5,$0$, and validation interval iterations 1000, so that we have about 50 checkpoints for every OCL model on every dataset.
-But to save storage, we only save the latter half of the total checkpoints.
-This setting is shared across all datasets.
-The batch size for image datasets is 32 for both training and validation, while for video datasets it is 8 for training and 4 for validation as there are more time steps in videos during validation.
-This setting is shared across all datasets.
-For multi-process, we use the number of workers 4. This holds for both training and validation. This setting is shared across all datasets.
-The objective is the CE and MSE loss following the basis methods original design.
-Metrics used here are ARI, mIoU and mBO, calculating the panoptic segmentation accuracy of both objects and the background. This setting is shared across all datasets.
-We use Adam optimizer with initial learning rate of 2e-4.
-And the learning rate is manipulated by cosine annealing scheduling, with a linear warmup during first 1/20 of total steps.
-This setting is shared across all datasets.
-We use auto mixed precision provided by PyTorch autocast API and compile.
-Along with this, we use PyTorch builtin gradient scaler, so as to apply gradient clipping of maximum norm 1.0 for images and 0.02 for videos.

### v5

@@
 Abstract.
-Perceiving visual scenes as objects and background—like humans do—Object-Centric Learning (OCL) aggregates image or video feature maps into object-level feature vectors, termed slots.
-OCL’s self-supervision of reconstructing the input from these aggregated slots struggles with complex object textures, thus Vision Foundation Model (VFM) representations are used as the aggregation input and reconstruction target.
-However, existing methods leverage VFM representations in diverse ways and often fail to fully exploit their potential.
-In response, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods.
-The key to our unification is simple yet effective, just shared quantizing the same VFM representation as the reconstruction target.
-Through mathematical modeling and statistical verification, we further analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision.
+Object-Centric Learning (OCL) aggregates image or video feature maps into object-level feature vectors, termed slots.
+It’s self-supervision of reconstructing the input from slots struggles with complex object textures, thus Vision Foundation Model (VFM) representations are used as the aggregation input and reconstruction target.
+Existing methods leverage VFM representations in diverse ways yet fail to fully exploit their potential.
+In response, we propose a unified architecture, Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO).
+The key to our unification is simply shared quantizing VFM representations in OCL aggregation and decoding.
 Experiments show that across different VFMs, aggregators and decoders, our VVO consistently outperforms baselines in object discovery and recognition, as well as downstream visual prediction and reasoning.
-The implementation and model checkpoints are available on https://github.com/Genera1Z/VQ-VFM-OCL.
-isbn: 978-1-4503-XXXX-X/2018/06submissionid: 7307copyright: none
+We also mathematically analyze why VFM representations facilitate OCL aggregation and why their shared quantization as reconstruction targets strengthens OCL supervision.
+Our source code and model checkpoints are available on https://github.com/Genera1Z/VQ-VFM-OCL.
+Object-Centric Learning, Vision Foundation Model, Vector Quantization, Object Representation, Visual Prediction, Visual Reasoning
+journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754566isbn: 979-8-4007-2035-2/2025/10submissionid: mfp7307copyright: noneccs: Computing methodologies Computer vision representations
 DINO2 ViT (s/14) ResNet18 (OCL fine-tune) distribution shift
@@
 Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects.
-Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024b).
+Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024).
 Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020).
@@
 OCL relies on pixel textures to discover objects.
-The early milestone (Locatello et al., 2020) reconstructs input pixels for supervision, which usually fails on realistic objects.
-Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, but at the cost of expensive annotations.
-Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, where super-pixels are codebook codes shared across samples, thus suppressing pixel redundancy and facilitating aggregation from the feature map into slots.
-Recent advances (Seitzer et al., 2023; Wu et al., 2023b) leverage Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract feature maps with better object separability, significantly improving OCL performance.
+The early milestone (Locatello et al., 2020) reconstructs input pixels as supervision, usually failing on realistic objects.
+Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, at the cost of expensive annotations.
+Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, whose super-pixels are codebook codes, thus suppressing pixel redundancy and facilitating aggregation from features into slots.
+Recent advances (Seitzer et al., 2023; Wu et al., 2023b) use Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract features with better object separability, boosting OCL significantly.
 However, existing OCL methods leverage VFM representations in quite different ways, as shown in Figure 3, and none of them fully utilize the power of VFM representations.
-To address these issues, we propose a clean architecture—Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)—that unifies mainstream OCL methods.
+To address these issues, we propose a clean architecture–Vector-Quantized VFMs for OCL (VQ-VFM-OCL, or VVO)–that unifies mainstream OCL methods.
 In this architecture, as shown in Figure 2, our VVO supports different VFMs for encoding, different OCL aggregators and different OCL decoders.
-The key to such unification is very simple—We quantize the representations from the same VFM, rather than from another encoder, as the reconstruction target.
-We also mathematically and statistically analyze why VFM representations facilitate OCL aggregation and how their shared quantization as reconstruction targets strengthens OCL supervision.
+The key to such unification is very simple–We quantize the representations from the same VFM, rather than from another encoder, as the reconstruction target.
+We also mathematically analyze why VFM representations facilitate OCL aggregation and why their shared quantization as reconstruction targets strengthens OCL supervision.
 Our contributions are: (i) A clean architecture, which unifies mainstream OCL methods.
@@
 2. Related Work
-SSseg vs OCL.
-Both Self-Supervised Segmentation (SSseg) (Ge et al., 2023; Wang et al., 2024a) and OCL can discovery objects, i.e., segmenting objects and the background in images or video frames without supervision.
-SSseg only outputs object segmentation masks, while OCL further represents objects as slots.
-These slots can be directly used in downstream object-centric world models (Wu et al., 2023a) for dynamics modeling, which is demanded in advanced vision tasks.
-Thus a direct comparison between them is beyond the scope of this work.
 OCL encoding.
-The stronger encoder that extracts feature maps of better objectness contributes more to OCL performance.
 Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder.
@@
 State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly.
-Although SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs, they remain unexploited in OCL setting.
+SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs yet remain unexploited in OCL setting.
 Our VVO supports various VFMs for OCL encoding.
@@
 3.1. Unify OCL
-As shown in Figure 2, our method adopts an architectural design of encoder-aggregator/quantizer-decoder.
-Firstly, OCL encoder $\bm{\phi}_{\mathrm{e}}$ transforms an input image or video frame $\bm{X}\in\mathbb{R}^{h_{0}\times w_{0}\times c_{0}}$, which is an observation of some visual scene, into a dense feature map $\bm{Z}\in\mathbb{R}^{h\times w\times c}$, for the following query-based aggregation:
+Our method adopts an architectural design as shown in Figure 2.
+Firstly, OCL encoder $\bm{\phi}_{\mathrm{e}}$ transforms input, an image or video frame $\bm{X}\in\mathbb{R}^{h_{0}\times w_{0}\times c_{0}}$ of some visual scene, into a dense feature map $\bm{Z}\in\mathbb{R}^{h\times w\times c}$, for the following query-based aggregation:
 (1) $\bm{\phi}_{\mathrm{e}}:\bm{X}\rightarrow\bm{Z}$
-where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained frozen VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024).
+where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024).
 As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here.
-We will detail the reasons in Section 3.2.
-Secondly, given queries $\bm{S}_{0}\in\mathbb{R}^{n\times c}$, OCL aggregator $\bm{\phi}_{\mathrm{a}}$ transforms $\bm{Z}$ into multiple feature vectors or slots $\bm{S}\in\mathbb{R}^{n\times c}$ and byproduct segmentation masks $\bm{M}\in\mathbb{R}^{h\times w}$, corresponding to objects or the background in the scene:
+We will explain it in Section 3.2.
+Secondly, given queries $\bm{S}_{0}\in\mathbb{R}^{n\times c}$, OCL aggregator $\bm{\phi}_{\mathrm{a}}$ transforms $\bm{Z}$ into multiple feature vectors or slots $\bm{S}\in\mathbb{R}^{n\times c}$ and corresponding byproduct segmentation masks $\bm{M}\in\mathbb{R}^{h\times w}$, each representing a specific object or background in the scene:
 (2) $\bm{\phi}_{\mathrm{a}}:(\bm{S}_{0},\bm{Z})\rightarrow(\bm{S},\bm{M})$
@@
 $\bm{M}$ is the binarized attention map thus intuitively reflects how well objects are represented by slots.
-For video OCL, there is a recurrent module turning current $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step. Such module can be Transformer encoder blocks (Vaswani et al., 2017).
+For video OCL, there is a recurrent module transitioning current slots $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step.
+Such module can be parameterized as a Transformer encoder block (Vaswani et al., 2017).
 Meanwhile, OCL quantizer $\bm{\phi}_{\mathrm{q}}$ transforms $\bm{X}$ into the reconstruction target $\bm{Q}\in\mathbb{R}^{h\times w\times c}$ for the following decoding:
 (3) $\bm{\phi}_{\mathrm{q}}:\bm{X}\rightarrow\bm{Q}$
-where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) module (Im Im et al., 2017; Van Den Oord et al., 2017).
+where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) (Im Im et al., 2017; Van Den Oord et al., 2017).
 But we meticulously design our own VQ variant, as detailed in Section 3.2.
@@
 We will explain why not use a separate typical VAE encoder in Section 3.2.
-Thirdly, OCL decoder $\bm{\phi}_{\mathrm{d}}$ transforms $\bm{S}$ into reconstruction $\bm{Q}^{\prime}\in\mathbb{R}^{h\times w\times c}$ with destructed $\bm{Q}$ as the condition:
+Thirdly, OCL decoder $\bm{\phi}_{\mathrm{d}}$ transforms slots $\bm{S}$ into reconstruction $\bm{Q}^{\prime}\in\mathbb{R}^{h\times w\times c}$ with destructed $\bm{Q}$ as the condition:
 (4) $\bm{\phi}_{\mathrm{d}}:(\bm{S},\mathrm{destruct}(\bm{Q}))\rightarrow\bm{Q}^{\prime}$
@@
 Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason.
-In Section 5.1, we analyze this both theoretically and empirically.
 Shared VFM Feature Quantization for Better Supervision
-Given the direct VFM feature map $\bm{Z}$, we transform it with a naive CNN to eliminate the positional information for better quantization.
-We quantize the adjusted $\bm{Z}$ using our VQ variant as $\bm{\phi}_{\mathrm{q}}$ to obtain the reconstruction target $\bm{Q}$.
+Given the VFM feature $\bm{Z}$, we adjust it via a small CNN to eliminate the positional information for better quantization; Then quantize it using our VQ variant $\bm{\phi}_{\mathrm{q}}$ as the reconstruction target $\bm{Q}$.
 Our VQ’s codebook follows SimVQ (Zhu et al., 2024).
@@
 (6) $\bm{T}:=\bm{W}\cdot\mathrm{sg}(\bm{T}_{0})$
-(7) $\bm{D}=||\bm{Z}-\bm{T}||_{2}^{2}$
-where $\mathrm{sg}(\cdot)$ is stop-gradient; $\bm{T}\in\mathbb{R}^{m\times c}$ is the codebook to be used for quantizing $\bm{Z}$; $\bm{D}\in\mathbb{R}^{h\times w\times m}$ is the matching distance between every super-pixel and every code.
-We convert distances to probabilities and select the most matched codes to form the quantization $\bm{Q}$:
+[…]
+where $\mathrm{sg}(\cdot)$ is stop-gradient; $\bm{T}\in\mathbb{R}^{m\times c}$ is the codebook for quantizing $\bm{Z}$; $\bm{D}\in\mathbb{R}^{h\times w\times m}$ is the matching distance between every super-pixel in $\bm{Z}$ and every code in $\bm{T}$.
+We convert distances to probabilities and select the most matched codes to form the quantization $\bm{Q}$ as the reconstruction target:
 (8) $\bm{P}=\mathrm{softmax}_{c}(-\bm{D})$
@@
 (11) $\bm{D}:=\frac{\bm{D}+\bm{G}}{\tau}$
-where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is the Gumbel noise and $\tau$ is the temperature.
-We devise annealing residual connection (Zhao et al., 2024a, b) after Equation 10 to stabilize pre-training:
+where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is Gumbel noise and $\tau$ is the temperature.
+Training-time annealing residual connection (Zhao et al., 2024a, b) is added after Equation 10 to stabilize pre-training:
 (12) $\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$
 where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing.
-Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$’s elements to be normalized:
-(13) $l_{n}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
+Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$ to be normal:
+(13) $l_{\mathrm{n}}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
 where $\lambda$ is empirically set to 0.1; $\mathbb{E}$ and $\mathbb{V}$ are calculated along height, width and channel dimensions.
-With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency and suppresses information redundancy or noise.
-This helps the aggregation with queries $\bm{S}_{0}$, which are also shared across samples.
-After such tokenization, our reconstruction target can be either regressed using mixture and diffusion decoding or classified using auto-regressive decoding.
-In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) use separate VAE and OCL encoders, which as shown in Figure 1 causes a distribution gap between $\bm{Q}$ and $\bm{Z}$.
+With all samples’ feature maps $\bm{Z}$ being represented with one codebook $\bm{T}$, the quantization $\bm{Q}$ naturally gains cross-sample consistency, helping the aggregation with queries $\bm{S}_{0}$, which are also shared across samples.
+Such tokenization is compatible with both regression and classification decoding.
+In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) are faced with distribution gaps between $\bm{Q}$ and $\bm{Z}$, shown in Figure 1, due to separate VAE and OCL encoders.
 Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision.
-In Section 5.2, we analyze this both theoretically and empirically.
-3.3. Compare Architectures
 Figure 3. Model architecture comparison.
-As shown in Figure 3, we compare the baselines with our VVO in a unified perspective. Specifically,
-- Our VVO: (1) VFMs are employed for OCL encoding and their features are fed to the aggregator directly, which eases OCL aggregation, as formulated in Section 3.2.
-(2) VFM features are shared quantized as reconstruction targets, which strengthens OCL supervision, as formulated in Section 3.2.
-- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator.
-(2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter.
-This loses much information of VFM features, while shared quantization can preserve such information (Van Den Oord et al., 2017).
-- SLATE (the unofficially improved version; we adopt it) (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b) / STEVE (Singh et al., 2022c): (1) Same as our VVO.
-(2) Reconstruction targets are discretized from features of a separate VAE encoder, instead of quantized from features shared from the same VFM for OCL encoding, which causes optimization noises.
-- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as our VVO.
-(2) Reconstruction targets are quantized from features of a separate VAE encoder, instead of sharing the same VFM for OCL encoding, which causes optimization noises.
-- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as our VVO.
-(2) Reconstruction targets are shared from the same VFM for OCL encoding but without quantization, which causes optimization noises.
-Our VVO adopts a clean and unified architecture with the above-mentioned two key designs.
-The first design, direct VFM feature extraction for OCL aggregation, is not proposed by us but we are the first to provide both theoretical and empirical analyses.
-The second design, shared VFM feature quantization as reconstruction targets, is proposed by us.
-Such shared quantization is also the key point that our VVO supports various OCL decoders.
-Specifically,
-- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features.
-As quantized features are actually the clustering centroids of continuous features, these decoders are inherently compatible with our shared quantized VFM features.
-- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are originally designed for discretized features, while (Seitzer et al., 2023) and (Kakogeorgiou et al., 2024) also explore their applicability to continuous features.
-Thus these decoders are naturally applicable to our shared quantized VFM features.
-- Diffusion-based OCl decoders, e.g., Conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), are originally designed for dicretized features.
-Thus if without our shared quantized VFM features as reconstruction targets, then this type of OCL decoders would not be unified into our architecture.
 Figure 4. Qualitative object discovery performance comparison.
 resolution= ClevrTex #slot=11 COCO #slot=7 VOC #slot=6 MOVi-D #slot=21 conditional
-256$\times$256 (224) ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
+$\times256$ (224) ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
 SLATE-DINO 17.4
@@
 resolution= ClevrTex #slot=11 COCO #slots=7 VOC #slot=6 MOVi-D #slot=21 conditional
-256$\times$256 ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
+$\times256$ ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
 SLATE-SAM 18.0
@@
 VVO is instantiated as VQSAM; Tfd, TfdT, Mlp and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
+3.3. Compare Architectures
+As shown in Figure 3, we compare baselines methods with our VVO in a unified perspective. Specifically,
+- Our VVO: (1) VFMs are employed for OCL encoding and their features are fed to the aggregator directly, which eases OCL aggregation, as formulated in Section 3.2.
+(2) VFM features are shared quantized as reconstruction targets, which strengthens OCL self supervision, as formulated in Section 3.2.
+- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator.
+(2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter.
+This loses much information of VFM features.
+- SLATE (the improved version; adopted here) / STEVE (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b): (1) Same as VVO.
+(2) Reconstruction targets are discretized from features of separate VAE encoding, not quantized from features shared from OCL encoding, causing optimization noises.
+- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as VVO.
+(2) Reconstruction targets are quantized from separate VAE encoding features, not sharing OCL encoding features, causing optimization noises.
+- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as VVO.
+(2) Reconstruction targets are shared from VFM features of OCL encoding without quantization, causing optimization noises.
+Our VVO realizes a clean and unified architecture based on the above-mentioned two key designs. Specifically,
+- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features, thus are compatible with our shared quantized VFM features.
+- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are designed for discretized features, while also showing applicability to continuous features (Seitzer et al., 2023; Kakogeorgiou et al., 2024).
+Thus they are applicable to our shared quantized VFM features.
+- Diffusion-based OCl decoders, e.g., conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), work on low-dimension features, necessitating our shared quantization on the continuous high-dimensional VFMs features.
 4. Experiment
@@
 contains synthetic videos, each with up to 20 daily objects dropping and bumping.
-COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its challenging panoptic segmentation.
+COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its instance segmentation.
 VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation.
-We also report results on real-world video dataset YTVIS version HQ [^6] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube.
+We also report results on real-world video dataset YTVIS [^6] https://youtube-vos.org/dataset/vis version HQ [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube.
 We choose Physion [^8] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
@@
 ARI ARIfg mBO mIoU
-Using higher resolution: 384$\times$384 (336)
-resolution=384$\times$384 (336) COCO #slot=7
+Using higher resolution: $\times384$ (336)
+[…]
 SLATE-DINO 41.4
@@
 Using different aggregators: SlotAttention, BO-QSA
-resolution=256$\times$256 (224) COCO #slot=7
+[…]
 SLATE-DINO-SlotAttention 17.0
@@
 Compared with general improvers: GDR and MSF
-resolution=256$\times$256 (224) COCO #slot=7
+[…]
 GDR
@@
 Compared with SotA methods: SPOT, VideoSAUR
-resolution=256$\times$256 (224) COCO #slot=7
+[…]
 SPOT-DINO 20.3
@@
 21.3 ±0.4 42.3 ±1.0 31.4 ±0.2 29.9 ±0.3
-resolution=256$\times$256 (224) YTVIS (HQ) #slot=7, unconditional
+[…]
 VideoSAUR-DINO 33.0
@@
 Table 4. VVO versus general improvers (upper) and SotA methods (lower) on object discovery. SPOT uses Transformer with 9 permutations Tfd9 as decoder while VideoSAUR uses SlotMixer SmdT.
-4.2. Evaluate on Object Discovery
-Object discovery task intuitively shows how well those slots separate different objects.
-We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^10] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_sco
-re.html
-(ARI), ARIfg (foreground), mean Intersection-over-Union [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
-With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO.
-As shown in Table 1, our method consistently improves object discovery performance across all types of OCL decoding.
-With a Transformer decoder Tfd for auto-regressive decoding, VVO significantly outperforms SLATE and STEVE across all datasets.
-With MLP Mlp for mixture decoding, VVO shows a smaller advantage over DINOSAUR but is still effective.
-With a conditional Diffusion model Dfz for diffusion decoding, VVO surpasses SlotDiffusion on most datasets.
-With supervised pretrained VFMs for OCL encoding, i.e., SAM2 Hiera+FPN (version t/16), our VVO is instantiated as VQSAM.
-As shown in Table 2, VVO consistently boosts those baseline methods’ object discovery performance across all decoding types on most datasets.
-As shown in Table 3, whether using a higher input resolution or different aggregators, VVO maintains its superiority over baselines.
-As shown in Table 4, VVO outperforms recent general improvers for OCL training, i.e., GDR and MSF, which do not support mixture decoding like us though.
-VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
-4.3. Evaluate on Set Prediction
 class labels bounding boxes
-top1$\uparrow$ R2$\uparrow$
+[…]
 DINOSAUR + MLP 0.61
@@
 0.64 ±0.1 0.59 ±0.1
-Table 5. Set prediction performance on COCO (panoptic, #slot=7).
+Table 5. Set prediction performance on COCO (#slot=7).
+4.2. Evaluate on Object Discovery
+Object discovery task intuitively shows how well those slots separate different objects.
+We evaluate all methods’ byproduct object segmentation accuracy with Adjusted Rand Index [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_sco
+re.html
+(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO) as metrics.
+With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO.
+As shown in Table 1, our method consistently improves object discovery performance across all types of OCL decoding.
+With a Transformer decoder Tfd for auto-regressive decoding, VVO significantly outperforms SLATE and STEVE across all datasets.
+With a spatial broadcast MLP decoder Mlp for mixture-based decoding, VVO shows a smaller advantage over DINOSAUR but is still effective.
+With a conditional Diffusion model Dfz for diffusion-based decoding, VVO surpasses SlotDiffusion on most datasets.
+With supervised pretrained VFMs for OCL encoding, i.e., SAM2 Hiera+FPN (version t/16), our VVO is instantiated as VQSAM.
+As shown in Table 2, VVO boosts all baselines’ object discovery performance across all decoding types on all datasets.
+As shown in Table 3, whether using higher input resolution or different aggregators, VVO maintains its superiority over baselines.
+As shown in Table 4, VVO outperforms recent OCL general improvers, i.e., GDR and MSF.
+VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
+4.3. Evaluate on Set Prediction
 Set prediction task directly shows how much object information those slots grasp.
 We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023).
-We measure top1 accuracy of the classified class labels while we measure the R2 score [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.r2_score.html of the regressed bounding box coordinates.
-As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO Mlp , obtains better object classification and bounding box regression.
-Thus, our models’ object representations are of higher quality.
+We measure top1 accuracy [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.accuracy_score.ht
+ml
+accuracy of the classified class labels while measure R2 score [^14] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.r2_score.html of the regressed bounding box coordinates.
+As shown in Table 5, compared with DINOSAUR, our VVO, i.e., VQDINO Mlp , obtains both better object classification and better object bounding box regression.
+Thus, our method extracts better slot representations for objects than the baseline.
 4.4. Deploy to the Downstream
@@
 We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO Mlp to extract slots.
-On visual prediction, we evaluate the prediction errors per time step measured in normalized MSE between regressed and extracted slots.
-As shown in Figure 5 upper, VVO accumulates prediction errors slower than the baseline.
-On visual reasoning, we evaluate the reasoning accuracies per time step between classified and ground-truth labels.
-As shown in Figure 5 lower, VVO accuracies are slightly lower at the beginning but much higher later than the baseline.
+On visual prediction, we evaluate the per time step prediction errors measured in normalized Mean Squared Error (MSE) between regressed and extracted slots.
+As shown in Figure 5 upper, our VVO accumulates prediction errors slower than the baseline.
+On visual reasoning, we evaluate the per time step reasoning accuracy between the classification outputs and ground-truth labels.
+As shown in Figure 5 lower, VVO’s accuracies are slightly lower at the beginning but much higher later than the baseline.
 4.5. Ablate the Architecture
 As shown in Table 6, VVO’s design of shared VAE and OCL encoder consistently outperforms separate VAE and OCL encoders, even when the latter employs another VFM for VAE encoding.
-Thus, VVO’s design of shared VFM representation quantization is crucial.
-However, the design of separate encoders remains prevalent among those baselines.
-As shown in Table 7, VVO’s improved quantizer variant, based on tricks of Gumbel noises defined in Equation 11, annealing residual connection defined in Equation 12 and normalizing regularization defined in Equation 13, is superior to the naive VQ.
+Thus, VVO’s design of shared VFM representation quantization is superior to the prevalent design of separate VAE and OCL encoders.
+As shown in Table 7, our improved quantizer variant for VVO, built upon tricks of Gumbel noises defined in Equation 11, annealing residual connection defined in Equation 12 and normalizing regularization defined in Equation 13, is superior to the naive VQ.
 The detailed effects of those tricks are shown in Figure 6.
@@
 VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders.
-Results are on COCO panoptic.
+Results are on COCO.
 shared VAE / Gumbel annealing normaliz. VQDINO
@@
 In comparison to VVO’s key designs, these tricks are more like the cherry on top.
-Results are on COCO panoptic with settings consistent with the above.
+Results are on COCO with settings consistent with the above.
 Figure 6.
@@
 Dfz
-on COCO panoptic.
+on COCO.
 5. Analysis
-We provide both mathematical modeling and statistical verification on why and how our VVO is better.
-5.1. Aggregation as Clustering
+We mathematically analyze our two key designs as below.
+Aggregation as Clustering
+As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).
+Denote the ideal query of $\bm{o}_{2}$ as $\bm{s}_{*}$, which is the clustering center or centroid of $\bm{o}_{2}$ and is closer to all super-pixels in $\bm{o}_{2}$ than in $\bm{o}_{1}$:
+(14) $d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$
+where $d(\cdot,\cdot)$ is a distance metric, e.g., minus inner product; $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary points in $\bm{o}_{1}$ and $\bm{o}_{2}$, respectively.
+But the actual query $\bm{s}$ follows $\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2}\bm{I})$. Substituting $\bm{s}$ for $\bm{s}_{*}$, the probability of correct aggregation is:
+(15) $p_{2}=p(d_{s1}>d_{s2})=\int_{\bm{v}\in\bm{o}_{2}}\frac{1}{\sqrt{2\pi}\bm{\sigma}}e^{-\frac{1}{2}(\frac{\bm{v}-\bm{s}_{*}}{\bm{\sigma}})^{2}}d\bm{v}$
+where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation hyper-plane between $\bm{o}_{1}$ and $\bm{o}_{2}$.
+The closer the boundary is to $\bm{s}_{*}$ the smaller the value of $p_{2}$ would be.
+According to Figure 1 the first observation, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space.
+This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space.
+Thus, $p_{2}$ in VFM space is bigger.
 Figure 7.
@@
 In VFM super-pixel space, the distance $d_{12}$ between $\bm{v}_{1}$ and $\bm{v}_{2}$ is larger, i.e., better objectness, thus $\bm{s}$ has higher probability $p(d_{s1}>d_{s2})$ to represent $\bm{o}_{2}$ correctly, compared with that in non-VFM super-pixel space.
-Mathematic Modeling
-As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).
-Denote the ideal query of $\bm{o}_{2}$ as $\bm{s}_{*}$, which is the clustering center or centroid of $\bm{o}_{2}$ and is closer to all super-pixels in $\bm{o}_{2}$ than in $\bm{o}_{1}$:
-(14) $d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$
-where $d(\cdot,\cdot)$ is a distance metric, e.g., minus inner product; $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary points in $\bm{o}_{1}$ and $\bm{o}_{2}$, respectively.
-But the actual query $\bm{s}$ follows $\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2}\bm{I})$. Substituting $\bm{s}$ for $\bm{s}_{*}$, the probability of correct aggregation is:
-[…]
-where $\bm{o}_{2}$ always contains $\bm{s}_{*}$, and is bounded by the separation plane between $\bm{o}_{1}$ and $\bm{o}_{2}$. The closer the boundary is to $\bm{s}_{*}$ the smaller $p_{2}$ is.
-According to Figure 1 the first observation, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space.
-This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space.
-Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space.
-$\square$
-Statistical Verification
-Equation 15 is very difficult to solve, but we can use Monte Carlo estimation:
-(16) $p_{2}=\frac{1}{N}\Sigma_{i=1}^{N}\mathbbm{1}[\bm{v}_{i}\in\bm{o}_{2}]$
-Specifically, we respectively use ResNet18, DINO2 ViT and SAM2 encoder to extract images into super-pixels, then apply KMeans clustering to find each clusters mean and std.
-From the Gaussian distributions defined by these mean and std, we repeatedly draw samples and calculate the ratio that such samples fall within their original clusters.
-This ratio is the correct aggregation probability.
-Figure 8. Statistical correlation between correct aggregation probability and object discovery accuracy on VOC.
-A step further, we can calculate the statistical correlation between the correct aggregation probability and the corresponding OCL performance.
-As shown in Figure 8, the correlation is roughly significant.
-Considering that correct aggregation probability can be calculated before the expensive OCL training, our Equation 15 is instructive to training better OCL models by selecting OCL encoders.
-5.2. Shared Quantization and Optimization Noise
-Mathematic Modeling
+Shared Quantization and Optimization Noise
 We reconstruct $\bm{Q}^{\prime}$ to approximate the target $\bm{Q}$ ultimately from $\bm{Z}$ via $\bm{\phi}_{\mathrm{a}}\circ\bm{\phi}_{\mathrm{d}}$, denoted as $f$ for simplicity.
@@
 Anyway, this has similar structure to Equation 16 and thus does not alter our judgment above.
-$\square$
-Statistical Verification
-Figure 9. Gradient distributions tend to disperse or have outliers if not using our shared VFM feature quantization. Results here are collected at time step 10k, on dataset VOC.
-The statistical bias in Equation 20 manifests as the gradient distribution of different layers, especially that of slots.
-No statistical bias is usually associated with a smooth zero-mean small-std Gaussian gradient distribution.
-Our technique, shared VFM feature quantization as reconstruction targets, includes two key points, sharing and quantization.
-Thus we compare three cases: (i) Shared and quantized, where DINO features are both used for aggregation and quantized for reconstruction; (ii) Shared but not quantized, where DINO features are used for aggregation and direct for reconstruction without quantization; (iii) Not shared but quantized, where DINO features are used for aggregation while quantized SAM features are used for reconstruction.
-As shown in Figure 9, using CE or MSE as the loss, our shared quantization achieves smoother Gaussian gradient distributions, which benefits OCL training.
-If not shared but quantized, then gradients tend to disperse; while if shared but not quantized, then gradients have some outliers.
 6. Conclusion
@@
 VVO has the potential to serve as a general testbed for research related to OCL in the future.
+Acknowledgements.
+We acknowledge the support of Finnish Center for Artificial Intelligence (FCAI), Research Council of Finland flagship program.
+We thank the Research Council of Finland for funding the projects ADEREHA (grant no.
+353198), BERMUDA (362407) and PROFI7 (352788).
+We also appreciate CSC - IT Center for Science, Finland, for granting access to supercomputers Mahti and Puhti, as well as LUMI, owned by the European High Performance Computing Joint Undertaking (EuroHPC JU) and hosted by CSC Finland in collaboration with the LUMI consortium.
+Furthermore, we acknowledge the computational resources provided by the Aalto Science-IT project through the Triton cluster.

### v6

@@
 But mainstream methods based on feature patches or a single feature vector disregard such compositionality.
-Inspired by human vision cognition (Bar, 2004; Cavanagh, 2011; Palmeri and Gauthier, 2004), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects.
-Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (Wu et al., 2023a), as well as in interactions between visual modality and other modalities (Zhang et al., 2024; Wang et al., 2024).
-Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (Locatello et al., 2020).
+Inspired by human vision cognition (bar2004visual; cavanagh2011visual; palmeri2004visual), Object-Centric Learning (OCL) decomposes visual scenes into multiple feature vectors, known as slots, each corresponding to an object or the background, thus enabling improved modeling of relationships and dynamics among objects.
+Object-centric representations have demonstrated superiority in advanced vision tasks, like prediction, reasoning, planning, and decision-making (wu2022slotformer), as well as in interactions between visual modality and other modalities (zhang2024omgllava; wang2024omnidrive).
+Existing OCL methods typically adopt an encoder-aggregator-decoder architecture (locatello2020slotattent).
 Firstly, the encoder transforms input image or video frame pixels into a dense feature map.
-Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (Locatello et al., 2020) with initial slots as the query.
+Then, the aggregator sparsifies this feature map into feature vectors via Slot Attention (locatello2020slotattent) with initial slots as the query.
 Lastly, the decoder reconstructs the input in some form from these aggregated slots, to provide the self-supervised training signal.
 OCL relies on pixel textures to discover objects.
-The early milestone (Locatello et al., 2020) reconstructs input pixels as supervision, usually failing on realistic objects.
-Some (Kipf et al., 2022; Elsayed et al., 2022) reconstruct optical flow or depth map to mitigate textural noises, at the cost of expensive annotations.
-Some (Singh et al., 2022a, c) reconstruct input’s VAE (Variational Autoencoder) representation, whose super-pixels are codebook codes, thus suppressing pixel redundancy and facilitating aggregation from features into slots.
-Recent advances (Seitzer et al., 2023; Wu et al., 2023b) use Vision Foundation Models (VFMs) (Caron et al., 2021; Oquab et al., 2023) to extract features with better object separability, boosting OCL significantly.
+The early milestone (locatello2020slotattent) reconstructs input pixels as supervision, usually failing on realistic objects.
+Some (kipf2021savi; elsayed2022savipp) reconstruct optical flow or depth map to mitigate textural noises, at the cost of expensive annotations.
+Some (singh2021slate; singh2022steve) reconstruct input’s VAE (Variational Autoencoder) representation, whose super-pixels are codebook codes, thus suppressing pixel redundancy and facilitating aggregation from features into slots.
+Recent advances (seitzer2023dinosaur; wu2023slotdiffuz) use Vision Foundation Models (VFMs) (caron2021dino1; oquab2023dino2) to extract features with better object separability, boosting OCL significantly.
 However, existing OCL methods leverage VFM representations in quite different ways, as shown in Figure 3, and none of them fully utilize the power of VFM representations.
@@
 OCL encoding.
-Early milestone methods like IODINE(Greff et al., 2019) and SA (Locatello et al., 2020) use small naive CNNs (Krizhevsky et al., 2012) as OCL encoder.
-Followups like SAVi (Kipf et al., 2022), SAVi++ (Elsayed et al., 2022), SLATE (Singh et al., 2022a) and STEVE (Singh et al., 2022c) employ pretrained ResNets (He et al., 2016), and fine-tune them on OCL datasets.
-State-of-the-arts like SlotDiffusion (Wu et al., 2023b) and DINOSAUR (Seitzer et al., 2023) utilize VFMs like DINO (Caron et al., 2021) and DINO2 (Oquab et al., 2023) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly.
-SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024) are also well recognized VFMs yet remain unexploited in OCL setting.
+Early milestone methods like IODINE(greff2019iodine) and SA (locatello2020slotattent) use small naive CNNs (krizhevsky2012alexnet) as OCL encoder.
+Followups like SAVi (kipf2021savi), SAVi++ (elsayed2022savipp), SLATE (singh2021slate) and STEVE (singh2022steve) employ pretrained ResNets (he2016resnet), and fine-tune them on OCL datasets.
+State-of-the-arts like SlotDiffusion (wu2023slotdiffuz) and DINOSAUR (seitzer2023dinosaur) utilize VFMs like DINO (caron2021dino1) and DINO2 (oquab2023dino2) ViTs (Vision Transformers) to extract highly object-separable feature map from input pixels, improving OCL performance significantly.
+SAM (kirillov2023sam) and SAM2 (ravi2024sam2) are also well recognized VFMs yet remain unexploited in OCL setting.
 Our VVO supports various VFMs for OCL encoding.
 OCL aggregation.
-SlotAttention (Locatello et al., 2020) is the footstone for mainstream OCL methods.
-Subsequent works like BO-QSA (Jia et al., 2023), ISA (Biza et al., 2023) and SysBind (Singh et al., 2022b) are all its variants, which are designed without changing the external interface.
+SlotAttention (locatello2020slotattent) is the footstone for mainstream OCL methods.
+Subsequent works like BO-QSA (jia2023boqsa), ISA (biza2023isa) and SysBind (singh2022sysbind) are all its variants, which are designed without changing the external interface.
 But considering their performance boosts, we only integrate BO-QSA by default.
@@
 With SlotAttention as the aggregator, the decoder and its reconstruction target affect OCL performance the most, as it is the source of supervision.
-Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (Zadaianchuk et al., 2024), decodes each slot’s spatial broadcast (Watters et al., 2019) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction.
-Transformer-based decoding, used in SLATE, STEVE and SPOT (Kakogeorgiou et al., 2024), reconstructs VAE representation of the input auto-regressively with slots as the condition.
-Diffusion-based decoding in LSD (Jiang et al., 2023) and SlotDiffusion drives slots to recover noise added to the input’s VAE representation.
+Mixture-based decoding, used in SAVi, SAVi++, DINOSAUR and VideoSAUR (zadaianchuk2024videosaur), decodes each slot’s spatial broadcast (watters2019spatialbroadcast) using naive CNNs or MLPs, and mixes them with corresponding weights into the reconstruction.
+Transformer-based decoding, used in SLATE, STEVE and SPOT (kakogeorgiou2024spot), reconstructs VAE representation of the input auto-regressively with slots as the condition.
+Diffusion-based decoding in LSD (jiang2023lsd) and SlotDiffusion drives slots to recover noise added to the input’s VAE representation.
 Our VVO supports all these types of OCL decoding.
 VAE for OCL.
-Variational Autoencoders (VAEs), like dVAE (Im Im et al., 2017) in SLATE and VQ-VAE (Van Den Oord et al., 2017) in SlotDiffusion, are employed to produce reconstruction targets for OCL training.
+Variational Autoencoders (VAEs), like dVAE (im2017dvae) in SLATE and VQ-VAE (van2017vqvae) in SlotDiffusion, are employed to produce reconstruction targets for OCL training.
 Since these VAEs are designed for image generation, some methods adapt them for OCL.
-Inspired by channel or weight grouping, GDR (Zhao et al., 2024a) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better.
-MSF (Zhao et al., 2024b) firstly exploits the multi-scale idea in the OCL setting with VAE-specific designs.
-Based on recent advancement RVQ (Yang et al., 2023) and SimVQ (Zhu et al., 2024), we design our own VQ variant for OCL.
+Inspired by channel or weight grouping, GDR (zhao2024gdr) decomposes features into attributes and combine them to produce VAE representation as reconstruction targets to guide OCL better.
+MSF (zhao2024msf) firstly exploits the multi-scale idea in the OCL setting with VAE-specific designs.
+Based on recent advancement RVQ (yang2023rvq) and SimVQ (zhu2024simvq), we design our own VQ variant for OCL.
 Figure 2.
@@
 (1) $\bm{\phi}_{\mathrm{e}}:\bm{X}\rightarrow\bm{Z}$
-where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained VFMs, like DINO (Caron et al., 2021), DINO2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) and SAM2 (Ravi et al., 2024).
+where $\bm{\phi}_{\mathrm{e}}$ can be parameterized as pretrained VFMs, like DINO (caron2021dino1), DINO2 (oquab2023dino2), SAM (kirillov2023sam) and SAM2 (ravi2024sam2).
 As OCL relies on textures to separate objects, $\bm{\phi}_{\mathrm{e}}$ should handle complex textures of objects, making VFMs necessary here.
@@
 (2) $\bm{\phi}_{\mathrm{a}}:(\bm{S}_{0},\bm{Z})\rightarrow(\bm{S},\bm{M})$
-where $\bm{\phi}_{\mathrm{a}}$ can be parameterized as widely adopted SlotAttention (Locatello et al., 2020) and its variants, which is some cross attention with $\bm{S}_{0}$ as queries and $\bm{Z}$ as keys and values.
+where $\bm{\phi}_{\mathrm{a}}$ can be parameterized as widely adopted SlotAttention (locatello2020slotattent) and its variants, which is some cross attention with $\bm{S}_{0}$ as queries and $\bm{Z}$ as keys and values.
 $\bm{M}$ is the binarized attention map thus intuitively reflects how well objects are represented by slots.
 For video OCL, there is a recurrent module transitioning current slots $\bm{S}$ into new queries $\bm{S}_{0}$ for the next time step.
-Such module can be parameterized as a Transformer encoder block (Vaswani et al., 2017).
+Such module can be parameterized as a Transformer encoder block (vaswani2017transformer).
 Meanwhile, OCL quantizer $\bm{\phi}_{\mathrm{q}}$ transforms $\bm{X}$ into the reconstruction target $\bm{Q}\in\mathbb{R}^{h\times w\times c}$ for the following decoding:
 (3) $\bm{\phi}_{\mathrm{q}}:\bm{X}\rightarrow\bm{Q}$
-where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) (Im Im et al., 2017; Van Den Oord et al., 2017).
+where $\bm{\phi}_{\mathrm{q}}$ can be parameterized as some Vector Quantization (VQ) (im2017dvae; van2017vqvae).
 But we meticulously design our own VQ variant, as detailed in Section 3.2.
@@
 (4) $\bm{\phi}_{\mathrm{d}}:(\bm{S},\mathrm{destruct}(\bm{Q}))\rightarrow\bm{Q}^{\prime}$
-Here $\bm{\phi}_{\mathrm{d}}$ can be parameterized as (i) a CNN or MLP for mixture decoding (Kipf et al., 2022; Seitzer et al., 2023), where $\bm{Q}$ is destructed to height and width, and $\bm{S}$ is spatially broadcast (Watters et al., 2019) into this shape and then decoded into components being mixed together; (ii) or a Transformer decoder for auto-regressive decoding (Singh et al., 2022a; Kakogeorgiou et al., 2024), where $\bm{Q}$ is destructed with causal masking as the query and $\bm{S}$ is the key and value; (iii) or a conditional Diffusion model for diffusion decoding (Wu et al., 2023b; Jiang et al., 2023), where $\bm{Q}$ is destructed with noise as the input and $\bm{S}$ is the condition.
+Here $\bm{\phi}_{\mathrm{d}}$ can be parameterized as (i) a CNN or MLP for mixture decoding (kipf2021savi; seitzer2023dinosaur), where $\bm{Q}$ is destructed to height and width, and $\bm{S}$ is spatially broadcast (watters2019spatialbroadcast) into this shape and then decoded into components being mixed together; (ii) or a Transformer decoder for auto-regressive decoding (singh2021slate; kakogeorgiou2024spot), where $\bm{Q}$ is destructed with causal masking as the query and $\bm{S}$ is the key and value; (iii) or a conditional Diffusion model for diffusion decoding (wu2023slotdiffuz; jiang2023lsd), where $\bm{Q}$ is destructed with noise as the input and $\bm{S}$ is the condition.
 Reconstructing $\bm{Q}$ using $\bm{S}$ drives $\bm{S}$ to aggregate as much object information as possible.
-Thus, a good reconstruction target, like in (Zhao et al., 2024a, b), is very important.
+Thus, a good reconstruction target, like in (zhao2024gdr; zhao2024msf), is very important.
 We will explain this in Section 3.2.
@@
 We directly extract the feature map $\bm{Z}$ from the input $\bm{X}$ using VFMs as $\bm{\phi}_{\mathrm{e}}$, where DINO2 ViT and SAM2 encoder are chosen and experimented thoroughly.
-No extra position encoding is needed here like in SA (Locatello et al., 2020) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$.
+No extra position encoding is needed here like in SA (locatello2020slotattent) because these VFMs already contain the positional information required in $\bm{\phi}_{\mathrm{a}}$.
 Since $\bm{\phi}_{\mathrm{e}}$ is frozen, We further use a trainable linear layer to adjust $\bm{Z}$ slightly.
 As shown in Figure 1, VFM representations have better objectness than non-VFMs, even under naive kMeans clustering [^1] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html.
-OCL aggregation is essentially a clustering whose initial centroids are trainable (Jia et al., 2023).
+OCL aggregation is essentially a clustering whose initial centroids are trainable (jia2023boqsa).
 Thus, we expect $\bm{\phi}_{\mathrm{a}}$ to aggregate VFM’s $\bm{Z}$ into slots $\bm{S}$ better under queries $\bm{S}_{0}$.
-Previous methods like DINOSAUR (Seitzer et al., 2023) have already exploited this but without any reason.
+Previous methods like DINOSAUR (seitzer2023dinosaur) have already exploited this but without any reason.
 Shared VFM Feature Quantization for Better Supervision
 Given the VFM feature $\bm{Z}$, we adjust it via a small CNN to eliminate the positional information for better quantization; Then quantize it using our VQ variant $\bm{\phi}_{\mathrm{q}}$ as the reconstruction target $\bm{Q}$.
-Our VQ’s codebook follows SimVQ (Zhu et al., 2024).
+Our VQ’s codebook follows SimVQ (zhu2024simvq).
 We predefine $m=4096$ template features, i.e., a codebook $\bm{T}_{0}\in\mathbb{R}^{m\times c_{0}}$, which are randomly initialized and remain frozen.
@@
 where $\mathrm{softmax}(\cdot)$ is calculated along the channel dimension; $\bm{P}$ is the match probabilities; $\bm{I}\in\mathbb{R}^{h\times w}$ is the matched code indexes; $\mathrm{argmax}(\cdot)$ is calculated along the channel dimension; $\mathrm{index}(\cdot,\cdot)$ is operated along code number dimension.
-The typical STE (Bengio et al., 2013) on $\bm{Q}$, needed in pre-training, can be skipped during OCL training.
+The typical STE (bengio2013ste) on $\bm{Q}$, needed in pre-training, can be skipped during OCL training.
 For its pre-training, we introduce some tricks. We add noise to $\bm{D}$ before Equation 8 to encourage code utilization:
@@
 where $\bm{G}\in\mathbb{R}^{h\times w\times m}$ is Gumbel noise and $\tau$ is the temperature.
-Training-time annealing residual connection (Zhao et al., 2024a, b) is added after Equation 10 to stabilize pre-training:
+Training-time annealing residual connection (zhao2024gdr; zhao2024msf) is added after Equation 10 to stabilize pre-training:
 (12) $\bm{Q}:=\alpha\bm{Z}+(1-\alpha)\bm{Q}$
 where $\alpha$ is scheduled from 1 to 0 during pre-training using cosine-annealing.
-Besides typical losses of reconstruction, alignment and commitment (Van Den Oord et al., 2017), we regularize the adjusted $\bm{Z}$ to be normal:
+Besides typical losses of reconstruction, alignment and commitment (van2017vqvae), we regularize the adjusted $\bm{Z}$ to be normal:
 (13) $l_{\mathrm{n}}=\lambda\mathrm{MSE}(\bm{Z},\mathrm{sg}(\frac{\bm{Z}-\mathbb{E}[\bm{Z}]}{\sqrt{\mathbb{V}[\bm{Z}]+\epsilon}}))$
@@
 Such tokenization is compatible with both regression and classification decoding.
-In contrast, methods like SLATE (Singh et al., 2022a) and SlotDiffusion (Wu et al., 2023b) are faced with distribution gaps between $\bm{Q}$ and $\bm{Z}$, shown in Figure 1, due to separate VAE and OCL encoders.
+In contrast, methods like SLATE (singh2021slate) and SlotDiffusion (wu2023slotdiffuz) are faced with distribution gaps between $\bm{Q}$ and $\bm{Z}$, shown in Figure 1, due to separate VAE and OCL encoders.
 Thus, we expect shared VFM representation quantization as reconstruction targets to strengthen OCL supervision.
@@
 (2) VFM features are shared quantized as reconstruction targets, which strengthens OCL self supervision, as formulated in Section 3.2.
-- SLATE0 (the official version) (Singh et al., 2022a): (1) VFM features are NOT directly fed to the aggregator.
+- SLATE0 (the official version) (singh2021slate): (1) VFM features are NOT directly fed to the aggregator.
 (2) VFM features are shared discretized to scalar numbers and re-embedded into features to be learned latter.
 This loses much information of VFM features.
-- SLATE (the improved version; adopted here) / STEVE (Jia et al., 2023; Singh et al., 2022c; Wu et al., 2023b): (1) Same as VVO.
+- SLATE (the improved version; adopted here) / STEVE (jia2023boqsa; singh2022steve; wu2023slotdiffuz): (1) Same as VVO.
 (2) Reconstruction targets are discretized from features of separate VAE encoding, not quantized from features shared from OCL encoding, causing optimization noises.
-- SlotDiffusion (Wu et al., 2023b) / LSD (Jiang et al., 2023): (1) Same as VVO.
+- SlotDiffusion (wu2023slotdiffuz) / LSD (jiang2023lsd): (1) Same as VVO.
 (2) Reconstruction targets are quantized from separate VAE encoding features, not sharing OCL encoding features, causing optimization noises.
-- DINOSAUR (Seitzer et al., 2023) / VideoSAUR (Zadaianchuk et al., 2024) / SPOT (Kakogeorgiou et al., 2024): (1) Same as VVO.
+- DINOSAUR (seitzer2023dinosaur) / VideoSAUR (zadaianchuk2024videosaur) / SPOT (kakogeorgiou2024spot): (1) Same as VVO.
 (2) Reconstruction targets are shared from VFM features of OCL encoding without quantization, causing optimization noises.
 Our VVO realizes a clean and unified architecture based on the above-mentioned two key designs. Specifically,
-- Mixture-based OCL decoders, e.g., CNN (Locatello et al., 2020; Kipf et al., 2022; Elsayed et al., 2022), MLP (Seitzer et al., 2023) and SlotMixer (Zadaianchuk et al., 2024), are originally designed for continuous features, thus are compatible with our shared quantized VFM features.
-- Auto-regressive OCL decoders, e.g., the Transformer decoder (Singh et al., 2022a, c) and Transformer9 (Kakogeorgiou et al., 2024), are designed for discretized features, while also showing applicability to continuous features (Seitzer et al., 2023; Kakogeorgiou et al., 2024).
+- Mixture-based OCL decoders, e.g., CNN (locatello2020slotattent; kipf2021savi; elsayed2022savipp), MLP (seitzer2023dinosaur) and SlotMixer (zadaianchuk2024videosaur), are originally designed for continuous features, thus are compatible with our shared quantized VFM features.
+- Auto-regressive OCL decoders, e.g., the Transformer decoder (singh2021slate; singh2022steve) and Transformer9 (kakogeorgiou2024spot), are designed for discretized features, while also showing applicability to continuous features (seitzer2023dinosaur; kakogeorgiou2024spot).
 Thus they are applicable to our shared quantized VFM features.
-- Diffusion-based OCl decoders, e.g., conditional Diffusion (Wu et al., 2023b; Jiang et al., 2023), work on low-dimension features, necessitating our shared quantization on the continuous high-dimensional VFMs features.
+- Diffusion-based OCl decoders, e.g., conditional Diffusion (wu2023slotdiffuz; jiang2023lsd), work on low-dimension features, necessitating our shared quantization on the continuous high-dimensional VFMs features.
 4. Experiment
@@
 We compare VVO with both OCL classics and state-of-the-arts.
-SLATE (Singh et al., 2022a) uses a Transformer decoder for auto-regressive decoding, and it differs from VVO in a separate VAE encoder and naive quantizer; STEVE (Singh et al., 2022c) is SLATE’s video version.
-DINOSAUR (Seitzer et al., 2023) uses an MLP for mixture decoding, and it differs from VVO in no quantization in its reconstruction target.
-SlotDiffusion (Wu et al., 2023b) uses a conditional Diffusion model for diffusion decoding, and it differs from VVO in a separate VAE encoder and naive quantizer.
-General improvers GDR (Zhao et al., 2024a) and MSF (Zhao et al., 2024b) only support auto-regression and diffusion decoding.
-We skip outdated methods like IODINE (Greff et al., 2019), SA (Locatello et al., 2020) and ISA (Biza et al., 2023) due to their low accuracy.
-We also skip SAVi (Kipf et al., 2022) and SAVi++ (Elsayed et al., 2022) as their extra modalities are unfair to others.
+SLATE (singh2021slate) uses a Transformer decoder for auto-regressive decoding, and it differs from VVO in a separate VAE encoder and naive quantizer; STEVE (singh2022steve) is SLATE’s video version.
+DINOSAUR (seitzer2023dinosaur) uses an MLP for mixture decoding, and it differs from VVO in no quantization in its reconstruction target.
+SlotDiffusion (wu2023slotdiffuz) uses a conditional Diffusion model for diffusion decoding, and it differs from VVO in a separate VAE encoder and naive quantizer.
+General improvers GDR (zhao2024gdr) and MSF (zhao2024msf) only support auto-regression and diffusion decoding.
+We skip outdated methods like IODINE (greff2019iodine), SA (locatello2020slotattent) and ISA (biza2023isa) due to their low accuracy.
+We also skip SAVi (kipf2021savi) and SAVi++ (elsayed2022savipp) as their extra modalities are unfair to others.
 Comparison.
@@
 re.html
-(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO) as metrics.
+(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (uijlings2013selectivesearch) (mBO) as metrics.
 With unsupervised pretrained VFMs for OCL encoding, i.e., DINO2 ViT (version s/14), our VVO is instantiated as VQDINO.
@@
 As shown in Table 4, VVO outperforms recent OCL general improvers, i.e., GDR and MSF.
-VVO also surpasses state-of-the-arts, i.e., SPOT (Kakogeorgiou et al., 2024) and VideoSAUR (Zadaianchuk et al., 2024), with their special types of decoding.
+VVO also surpasses state-of-the-arts, i.e., SPOT (kakogeorgiou2024spot) and VideoSAUR (zadaianchuk2024videosaur), with their special types of decoding.
 4.3. Evaluate on Set Prediction
 Set prediction task directly shows how much object information those slots grasp.
-We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023).
+We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (seitzer2023dinosaur).
 We measure top1 accuracy [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.accuracy_score.ht
@@
 We follow the convention to pretrain OCL models on Physion and represent this dataset as slots.
-Then the object-centric dynamics model SlotFormer (Wu et al., 2023a) is trained on those slots in an auto-regressive manner along the time dimension.
+Then the object-centric dynamics model SlotFormer (wu2022slotformer) is trained on those slots in an auto-regressive manner along the time dimension.
 We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO Mlp to extract slots.
@@
 Aggregation as Clustering
-As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (Locatello et al., 2020).
+As shown in Figure 7, super-pixels in a feature map $\bm{Z}$ all belong to two objects $\bm{o}_{1}$ and $\bm{o}_{2}$, so two queries are needed for aggregation, which is basically sum of super-pixels weighted by normalized minus distances between queries and super-pixels (locatello2020slotattent).
 Denote the ideal query of $\bm{o}_{2}$ as $\bm{s}_{*}$, which is the clustering center or centroid of $\bm{o}_{2}$ and is closer to all super-pixels in $\bm{o}_{2}$ than in $\bm{o}_{1}$: